### v2

# Low-Rank Interconnected Adaptation across Layers

Abstract
Low-rank adaptation (LoRA) is a powerful parameter-efficient fine-tuning method that utilizes low-rank projectors $A$ and $B$ to learn weight updates $\Delta W$ for adaptation targets $W$. Previous research has shown that LoRA is essentially a gradient compressor, performing random projections on the gradient using a fixed projection matrix $A_{0}$. However, this setup restricts the overall weight update to be low-rank, which limits the adaptation performance. In this paper, we propose low-rank interconnected adaptation across layers (Lily). Specifically, we employ a hierarchical framework where low-dimensional projectors (LPs) retained for downward projection at a particular level, while globally-shared high-dimensional projector (HP) experts perform upward projection across all levels of layers. Lily uniquely connects each LP to all HP experts, therefore the gradient projections are no longer dominated by fixed projection matrices, but rather by selective combinations of all the projectors, thereby breaking the low-rank constraint of LoRA. Furthermore, Lily’s cross-layer connections facilitate the capture of intricate information and dependencies across different layers, thereby enhancing the model’s representational capabilities. Experiments across various modalities, architectures, and model sizes underscore Lily’s great performance and efficiency. Code is available on github https://github.com/yibozhong/lily.
1 Introduction
For foundation models like Transformers (Vaswani et al., 2017b), fine-tuning on downstream tasks is a typical usage, but full fine-tuning (FFT) of large models like large language models (LLMs) incurs huge computational and storage costs and risks forgetting previously learned knowledge (Biderman et al., 2024). Linear probing, which fine-tunes only the final modules like classification heads, addresses these issues but leads to significant performance degradation since it doesn’t update weights from the backbone. To tackle these challenges, parameter-efficient fine-tuning (PEFT) has received significant attention. In PEFT, a model’s backbone weights are frozen, and lightweight trainable modules are introduced to efficiently learn task-specific knowledge. Among all PEFT methods, Low-rank Adaptation (LoRA (Hu et al., 2021)) is one of the most widely applied techniques, especially in LLMs. LoRA introduces a pair of low-rank projection matrices for each adaptation target, consisting of a downward adapter $A$ and an upward adapter $B$, to approximate $\Delta W$ in FFT. Due to its low-rank nature, LoRA offers significant computational and storage savings, effectively alleviating the burdens of FFT while significantly outperforming linear probing by learning the weight updates for backbone weight.
However, LoRA and many subsequent improvements to the method (Miles et al., 2024), (Zhang et al., 2023), (Zhong et al., 2024) have a limitation: the overall learned weight updates $\Delta W$ are restricted to be low-rank, which limits the model performance during adaptation. Previous work (Hao et al., 2024) has shown that LoRA adapters are secretly gradient compressors. In particular, LoRA updates to the weight can be approximated by performing random projection on the gradient, using a fixed random projection matrix $A_{0}$, where $A_{0}$ is the initial weight for $A$. Since this limitation arises because the projection matrix is always fixed in LoRA, using multiple distinct matrices and their variations to perform the gradient projection could avoid restricting the weight updates to a low-rank subspace. This prompts a question: How can we enable high-rank updates by using various projection matrices for an adaptation target?
Figure 1: Dynamics of LoRA and Lily. $N_{l}$ is used to denote the number of layers in the model while $N_{e}$ denoting the number of high-dimensional projector experts from Lily. R is representing the router from Lily.
In this paper, we propose Low-rank interconnected adaptation across layers (Lily), a novel framework for more expressive and performative PEFT. Specifically, we decouple the downward low-dimensional projector (LP) and its corresponding upward high-dimensional projectors (HP), making them not tightly-bonded. Each LP is connected to all the HPs, and vice versa as shown in Fig. 1. This results in a hierarchical structure where LPs are still retained at a particular level to perform downward projection, while all HPs are now globally shared by all the LPs, performing upward projection. Inspired by self-attention Vaswani et al. (2017a), which calculates the relationship between a token and all tokens and obtains attention scores indicating the strength of their relationship, we selectively connect an LP with the HPs based on layer features. The LP extracts features from the current layer, and based on the extracted features, a data-dependent and selective combination of HPs is performed. This is realized by utilizing a router (Shazeer et al., 2017) that outputs a unique weight distribution for HP experts, depending on the current input feature, thereby exhibiting selectivity.
The gradient projection now uses multiple random projection matrices from all the layers and their selective combinations rather than a fixed matrix, thus effectively avoiding the constraint of low-rank updates from LoRA. Furthermore, Lily enables a more comprehensive information access by allowing adapters at each layer to access information from other layers, promoting an interconnected and dynamic learning process, where the adapters can collaborate, share learned knowledge and model dependencies across layers. Overall, our contributions include:
  • We propose Lily, a novel PEFT framework that incorporates cross-layer connections of the projection matrices, thereby breaking the restriction of low-rank weight updates in LoRA.
  • Lily utilizes routers to selectively connect an LP with multiple HP experts, enabling comprehensive information access and therefore expressive adaptation.
  • Extensive experiments are conducted across various modalities, architectures, and model sizes, highlighting Lily’s great performance and efficiency in diverse scenarios.
2 Related Work
Parameter Efficient Fine-Tuning. Typical usage of foundation models includes pre-training on large datasets and fine-tuning on various downstream tasks. Parameter-efficient fine-tuning (PEFT) thus emerges as a promising field, aiming to fine-tuning the model efficiently with minimal parameters while maintaining performance and preserving previously learned knowledge, addressing drawbacks posed by conventional fine-tuning techniques like full fine-tuning or linear probing. Current PEFT research can be mainly categorized into two types: 1) adapter-based methods (Hu et al., 2021), (Chen et al., 2022), (Pfeiffer et al., 2020b), (Jie and Deng, 2023) (Houlsby et al., 2019b) and 2) prompt-based methods (Tu et al., 2023b) (Tu et al., 2023a). Adapter-based methods introduce lightweight adapters into the Multi-Head Self-Attention (MHSA) or the Feed-Forward Network (FFN) blocks within the Transformer architecture. On the other hand, prompt-based methods append trainable tokens as prompts to the input sequence fed to certain parts of the model.
Among these various PEFT techniques, low-rank adaptation (LoRA (Hu et al., 2021)) stands out as one of the most well-known methods. LoRA introduces a pair of projection matrices $A$ and $B$ per adaptation target $W$. The low-dimension projector (LP) $A$ projects input $x$ to low-dimension space, and the high-dimension projector (HP) $B$ restores it to its original dimension. Multiplying these projection matrices approximates the weight update $\Delta W$ in FFT. Recent work (Hao et al., 2024) has shown that LoRA adapters are essentially performing random projection to the gradient using a fixed matrix. This restricts the learned weight update to low-rank subspace and thus imitating the model performance. Meanwhile, $A$ and $B$ are tightly coupled, therefore the adaptation process only has information access from current layer, without an understanding of information from other layers, which could be beneficial to modeling dependencies across various layers.
Mixture of Experts. Mixture of Experts (MoE) is an active research area that has garnered significant attention, especially in the field of large language models (LLMs). Conditional computation, where different parts of the network are activated on a per-example basis, has been proposed to enhance model capability without increasing computation (Davis and Arel, 2013) (Bengio et al., 2013) (Eigen et al., 2013) (Almahairi et al., 2016). The sparsely-gated MoE layer is introduced to implement this idea, consisting of numerous sub-networks (Shazeer et al., 2017). A trainable gating network, known as a "router", determines the combination of experts for each example. There are already PEFT methods like MoLORA (Zadouri et al., 2023) and MOLA (Gao et al., 2024a) which apply the MoE design concept to PEFT. However, these methods simply treat the adapters combined in LoRA as a single expert. A concurrent research Wu et al. (2024), utilizes LP and HP sub-spaces as the experts but fails to overcome the limitation discussed in previous section.
3 Methodology
3.1 Preliminaries
LoRA is a Gradient Projector. Adaptation process of low-rank adaptation (LoRA) can be expressed as follows when setting the rank to $d$, a small hidden-dimension:
$W=W_{0}+AB$ (1)
where $W,W_{0}\in\mathbb{R}^{C_{in}\times C_{out}}$ with $C_{in}$ being the input dimension and $C_{out}$ being the output dimension. $W_{0}$ is the frozen weight of a target module and $W$ is the actual value of the target during adaptation runtime. $A\in\mathbb{R}^{C_{in}\times d}$ and $B\in\mathbb{R}^{d\times C_{out}}$ are low-dimension projectors (LP) and high-dimension projectors (HP), with their product multiplied by a scalar value $s$ (omitted in the equation) to approximate weight update $\Delta W\in\mathbb{R}^{C_{in}\times C_{out}}$. We denote the information from $W_{0}$ as the basic knowledge that the pre-trained model already possesses, while $\Delta W=AB$ represents the learned extra knowledge specific to the current downstream task. We denote the initial value of $A$ and $B$ as $A_{0}$ and $B_{0}$, and their update as $\Delta A$ and $\Delta B$. $B_{0}$ is initialized as 0. The approximation is then described as:
$W$ $=W_{0}+(A_{0}+\Delta A)(B_{0}+\Delta B)$ (2)
$=W_{0}+A_{0}\Delta B+\Delta A\Delta B$ (3)
$\approx W_{0}+A_{0}\Delta B$ (4)
It is proved in Hao et al. (2024) that it is the term $A_{0}\Delta B$ that contributes the most to the weight updates. The forward pass given an input in LoRA is:
$y=xW=xW_{0}+xAB$ (5)
where $x\in\mathbb{R}^{C_{in}}$ and $y\in\mathbb{R}^{C_{out}}$ for simplicity. Therefore, the gradient w.r.t. $B$ is:
$\frac{\partial{\mathcal{L}}}{\partial B}$ $=A^{T}x^{T}\frac{\partial{\mathcal{L}}}{\partial y}$ (6)
$=A^{T}(\nabla_{W}{\mathcal{L}})$ (7)
where $x^{T}\frac{\partial{\mathcal{L}}}{\partial y}\in\mathbb{R}^{C_{in}\times C_{out}}$ is the gradient for the weight matrix, which we denote as $(\nabla_{W}{\mathcal{L}})$. Using the approximation in Eq. 2, we can express the adaptation when only updating $B$ using gradient descent:
$W=W_{0}-\eta\sum_{t=0}^{T}\big{[}A_{0}A_{0}^{T}(\nabla_{W}{\mathcal{L}}_{t})\big{]}$ (8)
where $T$ is the current time-step and $\eta$ is the adopted learning rate. The projection matrix here is $A_{0}$ because $A$ is not updated in the approximation, since its change contribute little to the overall weight updates. We can observe that LoRA is essentially performing random projection to the gradient using a fixed random projection matrix $A_{0}$. However, this setup limit the weight update $\Delta W$ to be low-rank, which also limit the adaptation performance.
Lily is a Better Gradient Projector. Consider simply connecting each $A$ with all the $B$s with weights, the forward pass is then:
$W$ $=W_{0}+\sum_{i=1}^{N_{e}}\Big{[}S_{i}\big{[}(A_{0}+\Delta A)(B_{0}^{i}+\Delta B^{i})\big{]}\Big{]}$ (9)
$=W_{0}+\sum_{i=1}^{N_{e}}\big{[}S_{i}(A_{0}\Delta B^{i}+\Delta A\Delta B^{i})\big{]}$ (10)
$\approx W_{0}+A_{0}\sum_{i=1}^{N_{e}}S_{i}\Delta B^{i}$ (11)
where $N_{e}$ is used to denote the number of $B$s in the model and $S$ is a set of weight scores of current $A$ for all the $B$s, each $S_{i}$ is a scalar value representing the weight for corresponding $B^{i}$.
Thanks to the inter-connectivity, each $B$ is also connected to all the $A$s. The gradient w.r.t an $B$ is now:
$\frac{\partial{\mathcal{L}}}{\partial B}$ $=\sum_{i=1}^{N_{l}}C_{i}{A^{i}}^{T}{x^{i}}^{T}\frac{\partial{\mathcal{L}}}{\partial y^{i}}$ (12)
$=\sum_{i=1}^{N_{l}}C_{i}{A^{i}}^{T}(\nabla_{W_{i}}{\mathcal{L}})$ (13)
where $N_{l}$ is the number of $A$s in the model and we denote gradient for each corresponding weight matrix as $\nabla_{W_{i}}{\mathcal{L}}$. $C$ is a set of weight scores where each $C_{i}$ is a weight score from certain $A_{i}$ for current $B$. The forward pass when only updating $B$ can then be expressed as:
$W=W_{0}-\eta\sum_{t=0}^{T}\Big{[}\sum_{1}^{N_{e}}S_{i}\big{[}\sum_{j=1}^{N_{l}}{C_{i,j}A_{0}A_{0}^{j}}^{T}(\nabla_{W_{j}}{\mathcal{L}}_{t})\big{]}\Big{]}$ (14)
where $C_{i,j}$ is used to denote weight scores from various $A^{j},j\in\{1,2,...,N_{l}\}$ to $B_{i},i\in\{1,2,...,N_{e}\}$. It can be observed that with this setup, the projection to the gradient no longer use a fixed $A_{0}$, but rather utilizing all the projection matrix $A_{0}^{i},i\in\{1,2,...,N_{l}\}$ and their overall complex combinations. Note that all the ${A_{0}}^{i}$s are distinct random projection matrices and each weight score ($S_{i}$ and $C_{i,j}$) is unique, data-dependent and thus varies across time-steps. Therefore, the weight updates are not restricted to be low-rank, but rather high-rank. This setup is precisely what Lily adopts. To differentiate from LoRA, we use different notations for the projectors. Specifically, we decouple HPs from their original layers, allowing them to form a model-wide shared module of HP experts ($B$ in LoRA), denoted as $P_{H}^{i}$, where $i\in\{1,2,...,N_{e}\}$ and $N_{e}$ is the number of HP experts. Meanwhile, we fix LPs ($A$ in LoRA) $P_{L}^{i},i\in\{1,2,...,N_{l}\}$ in each layer for downward projection, where $N_{l}$ is the number of LPs. This hierarchical structure makes the projection matrices all interconnected across layers, enabling the modeling of complex dependencies and high-rank updates to the model weights. Detailed model structure and design intuitions are discussed in Appendix A.
3.2 Dynamics of Lily
Downward Projection and Selective Weight Allocation. The process is illustrated in the right half of Fig. 1. Initially, we use an LP to project the input $x\in\mathbb{R}^{N\times C_{in}}$ into its low-dimensional representation $x^{\prime}\in\mathbb{R}^{N\times d}$ where $N$ is the sequence length:
$x^{\prime}=xP_{L}$ (15)
Inspired by the Mixture of Experts (MoE) paradigm, we employ a router $R\in\mathbb{R}^{N_{e}\times d}$ to selectively assign weights to all HP experts based on their relationship to the current layer’s features ($x^{\prime}$). The weight set $S$ is obtained as:
$S=softmax(\sum_{i=1}^{N}(x^{\prime}R^{T})_{i})$ (16)
The router selectively combines experts based on the current layer’s features, enabling smart information integration. For shallower inputs, the router increases attention for experts specializing in shallow-layer knowledge, while deeper inputs favor experts learning deep-layer knowledge.
Weighted Combination of Experts and Upward Projection. Once we obtain the low-dimensional input $x^{\prime}$ , we combine information from all layers using the model-wide shared global HP module. One intuitive approach is to feed $x^{\prime}$ into each HP expert and combine their outputs to obtain the extra knowledge $x_{\Delta}\in\mathbb{R}^{N\times C_{out}}$:
$x_{\Delta}=\sum_{i=1}^{N_{e}}S_{i}\cdot(x^{\prime}P_{H}^{i})$ (17)
where $S\in\mathbb{R}^{N_{e}}$ is the set of weight scores for HP experts, obtained through selective weight allocation. To address efficiency concerns discussed in Appendix A.2, we propose an alternative implementation that is mathematically equivalent and significantly reduces computational burden, described as:
$x_{\Delta}=x^{\prime}(\sum_{i=1}^{N_{e}}S_{i}\cdot P_{H}^{i})$ (18)
Since each $S_{i}$ is a scalar value, the calculation in Eq. 18 is mathematically equivalent to that in Eq. 17, but with significantly improved efficiency. Therefore, the whole computation flow, with input $x\in\mathbb{R}^{N\times C_{in}}$ and output $y\in\mathbb{R}^{N\times C_{out}}$, for an adaptation target module is:
$y$ $=xW_{0}+s\cdot x_{\Delta}$ (19)
$=xW_{0}+s\cdot xP_{L}(\sum_{i=1}^{N_{e}}(softmax(\sum_{j=1}^{N}(xP_{L}R^{T})_{j}))_{i}\cdot P_{H}^{i})$ (20)
By selectively allocating weights and combining HP experts, Lily enables access to all levels of information during adaptation. Each layer’s target adaptation modules could consider the status and knowledge from all other layers, resulting in a more expressive and comprehensive adaptation. Meanwhile, thanks to its inter-connectivity and selectivity, Lily break the low-rank update constraint of LoRA and enable high-rank updates, as discussed in preliminaries.
4 Experiments
We validate the effectiveness of Lily across different domains, model sizes (from ViT to LLM), and architectures (Transformers, Mamba), demonstrating its general strong adaptation capability. Concurrently, we conduct a comprehensive analysis of Lily’s intrinsic mechanisms, providing a thorough understanding of Lily. All experiments are conducted on a single RTX 4090 GPU. Additionally, multiple analysis are provided in Appendix C, D, E, F, G, H, I and J.
4.1 Common Sense Reasoning
Table 1: Commonsense reasoning results for Falcon-Mamba-7B across eight tasks. Bold represents the highest performance for each dataset utilizing PEFT methods.
Model PEFT BoolQ PIQA SIQA HellaSwag WinoGrande ARC-e ARC-c OBQA Avg.
ChatGPT - 73.1 85.4 68.5 78.5 66.1 89.8 79.9 74.8 77.0
Falcon-Mamba-7B LoRA 6.5 30.5 40.6 14.9 56.4 42.2 31.8 38.4 32.7
Lily (Δ + $\Delta\text{ + in}$) 44.9 66.8 65.0 10.5 57.1 78.7 64.6 68.2 57.0
Lily (in) 60.2 61.0 67.3 12.9 61.5 80.0 67.5 65.8 59.5
Table 2: Commonsense reasoning results for LLaMA3-8B across eight tasks. represents results taken from Liu et al. (2024). Bold denotes the highest performance scores for each dataset among different PEFT methods.
Model PEFT BoolQ PIQA SIQA HellaSwag WinoGrande ARC-e ARC-c OBQA Avg.
ChatGPT - 73.1 85.4 68.5 78.5 66.1 89.8 79.9 74.8 77.0
LLaMA3-8B LoRA 70.8 85.2 79.9 91.7 84.3 84.2 71.2 79.0 80.8
PiSSA 67.1 81.1 77.2 83.6 78.9 77.7 63.2 74.6 75.4
MiLoRA 68.8 86.7 77.2 92.9 85.6 86.8 75.5 81.8 81.9
Lily 72.9 85.6 77.8 92.7 83.3 89.7 77.6 82.8 82.8
Implementation We evaluate Lily on commonsense reasoning with LLMs. Regarding the implementation, we utilize LLaMA3-8B (AI@Meta, 2024) and Falcon-Mamba-7B (Zuo et al., 2024) as backbones. LLaMA3 is a near-SOTA open-source large language model, while Falcon-Mamba is the latest and only open-source large language model based on the Mamba architecture. Using these models allows us to validate the effectiveness of Lily for fine-tuning LLMs and whether this effectiveness can be transferred to architectures beyond Transformers (Mamba, in this case). We fine-tune these models on Commonsense170K (Hu et al., 2023) and evaluate the adaptation results on eight multiple-choice problem tasks, including BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-e, ARC-c (Clark et al., 2018), and OBQA (Mihaylov et al., 2018). The compared methods are LoRA for Falcon-Mamba and LoRA (Hu et al., 2021), PiSSA (Meng et al., 2024), and MiLoRA (Wang et al., 2024) for LLaMA3. We only compare LoRA for Falcon-Mamba because tailored PEFT methods for Mamba-based LLMs have not yet been proposed, which is beyond the scope of this paper. Detailed hyper-parameter settings and datsets information are reported in Appendix B.1.1 and Appendix B.2.1.
Results We report the accuracy in the Table 2 and Table 1. Based on the results, it is evident that Lily performs the best out of the compared PEFT methods. Lily surpasses LoRA by a significant margin on Falcon-Mamba, and on LLaMA3, it outperforms LoRA and MiLoRA. This indicates Lily’s superior adaptation capability and parameter efficiency dealing with commonsense reasoning tasks. Additionally, while the performance on Falcon-Mamba is notably lower than the baseline and LLaMA3, we believe this is due to the model’s limitations rather than Lily’s, as Lily still significantly outperforms LoRA on Falcon-Mamba and demonstrates great performance on LLaMA3. This sheds light on the current state of Mamba-based LLMs, showing that they generally have inferior performance compared to Transformer-based LLMs like ChatGPT and LLaMA on many tasks.
4.2 Natural Language Understanding
Table 3: Various fine-tuning methods applied to RoBERTa Base are evaluated on 6 datasets from the GLUE benchmark. We present the Matthew’s correlation coefficient (MCC) for CoLA, Pearson correlation coefficient (PCC) for STS-B, and accuracy (Acc.) for the remaining tasks. The highest performance for each dataset is highlighted in bold, with all metrics favoring higher values across the 6 datasets.
Model & Method
# Trainable
Parameters
SST-2
MRPC
CoLA
QNLI
RTE
STS-B
Avg.
$\rm{RoB_{base}}$(FFT) 125M 94.8 90.2 63.6 92.8 78.7 91.2 85.2
$\rm{RoB_{base}}$(BitFit) 0.1M 93.7 92.7 62 91.8 81.5 90.8 85.4
$\rm{RoB_{base}}$($\text{Adpt}^{\text{D}}$) 0.3M 94.2 88.5 60.8 93.1 71.5 89.7 83.0
$\rm{RoB_{base}}$($\text{Adpt}^{\text{D}}$) 0.9M 94.7 88.4 62.6 93.0 75.9 90.3 84.2
$\rm{RoB_{base}}$(LoRA) 0.3M 95.1 89.7 63.4 93.3 78.4 91.5 85.2
$\rm{RoB_{base}}$(AdaLoRA) 0.3M 94.5 88.7 62.0 93.1 81.0 90.5 85.0
$\rm{RoB_{base}}$(DyLoRA) 0.3M 94.3 89.5 61.1 92.2 78.7 91.1 84.5
$\rm{RoB_{base}}$(Lily) 0.3M 95.0 90.2 66.0 92.5 81.6 90.8 86.0
Implementation We evaluate Lily on natural language understanding (NLU) tasks. For implementation, we use RoBERTa Base (Liu et al., 2019) as the backbone and fine-tune it on tasks from GLUE benchmark (General Language Understanding Evaluation (Wang et al., 2018)), consisting of multiple NLU tasks including single-sentence classification tasks, similarity and paraphrase tasks and natural language inference tasks. We compare Lily against several competitive PEFT methods, including BitFit (Zaken et al., 2021), Adapter-Tuning (Rücklé et al., 2020) (Houlsby et al., 2019a) (Lin et al., 2020) (Pfeiffer et al., 2020a), LoRA (Hu et al., 2021), DyLoRA (Valipour et al., 2022) and AdaLoRA (Zhang et al., 2023). Additionally, we utilize full fine-tuning (FFT) as the baseline. Specific hyper-parameters and datasets information are provide in Appendix B.1.2 and B.2.2.
Results The results are shown in Table. 3, from which we can clearly observe that Lily surpass all of the compared PEFT methods by a significant margin, demonstrating its capability of tackling NLU tasks. Among the 6 given tasks, Lily surpasses FFT in 4 of them, showcasing its strong approximation ability with compatible parameter-efficiency.
Figure 2: Results of subject-driven generation. Lily’s results align better with prompts, featuring more accurate color, environment, and shape.
4.3 Subject-driven Image Generation
Implementation We conduct experiments on fine-tuning text-to-image diffusion models for the subject-driven generation task (Ruiz et al., 2023). For backbone, we use SDXL and we fine-tune it using LoRA and Lily. We first fine-tune the model with images associated with text prompts (e.g., A photo of a [v] duck toy), in which a unique identifier is provided. After that, text prompts containing the identifier could be used to generate customized images.
Results The results are presented in Fig. 2 following the format in Gao et al. (2024b) and Wu et al. (2024), from which we can observe that images generated by Lily generally align better with the text prompts. For instance, when asked to generate a duck toy floating on top of water, Lily’s image accurately depicts the designated environment, whereas LoRA’s does not. Additionally, when asked to generate a wolf plushie in snow, Lily precisely depicts the snow around the wolf, while LoRA fails to do so. These observations demonstrate Lily’s excellent ability in the domain of text-to-image generation with more expressive adaptation. More generated results are in Appendix I.
4.4 Visual Adaptation Benchmark
Table 4: Full results of Lily on ViT-B pre-trained on ImageNet-21K for the VTAB-1K benchmark, with averages computed based on group-wise results. Bold indicates the best performance.
Natural Specialized Structured
Params(M)
Average
Cifar100
Caltech101
DTD
Flowers102
Pets
SVHN
Sun397
Camelyon
EuroSAT
Resisc45
Retinopathy
Clevr-Count
Clevr-Dist
DMLab
KITTI-Dist
dSpr-Loc
dSpr-Ori
sNORB-Azim
sNORB-Ele
Conventional Fine-Tuning
FFT 86 68.9 68.9 87.7 64.3 97.2 86.9 87.4 38.8 79.7 95.7 84.2 73.9 56.3 58.6 41.7 65.5 57.5 46.7 25.7 29.1
LP 0 57.6 64.4 85.0 63.2 97.0 86.3 36.6 51.0 78.5 87.5 68.5 74.0 34.3 30.6 33.2 55.4 12.5 20.0 9.6 19.2
PEFT methods
AdaptFormer 0.588 76.8 74.0 92.2 71.7 99.3 91.7 88.9 56.4 87.2 95.1 85.7 75.9 84.2 62.2 53.0 81.0 87.1 53.6 35.3 42.3
Bi-LoRA 1.180 76.7 72.1 91.7 71.2 99.1 91.4 90.2 55.8 87.0 95.4 85.5 75.5 83.1 64.1 52.2 81.3 86.4 53.5 36.7 44.4
LoRA 1.180 76.4 72.5 91.5 71.9 99.1 91.4 89.6 56.0 87.6 95.3 84.0 75.0 83.6 64.3 51.6 80.9 86.0 51.8 36.8 42.3
FourierFT 0.936 72.7 69.1 88.8 71.9 99.0 91.0 79.0 55.6 84.9 93.0 83.2 74.9 70.7 61.1 45.2 74.8 78.0 53.0 24.8 30.8
MoRA 1.058 75.4 72.1 90.0 71.7 99.2 91.1 90.1 56.0 87.1 94.8 85.1 75.4 76.7 62.3 49.7 78.3 83.1 53.0 34.5 34.5
Lily 0.318 77.3 73.9 93.0 72.9 99.3 91.6 89.0 56.6 87.9 95.2 84.9 75.7 83.9 65.4 53.4 81.6 88.2 54.5 37.0 45.4
Table 5: Full results of Lily on Vim-S pre-trained on ImageNet-1K for the VTAB-1K benchmark, with averages calculated within each group. * denotes linear probing results from Tu et al. (2023b). For fair comparison, we also use ViT-B pre-trained on ImageNet-1K. Bold indicates best performance among Vim-based PEFT methods.
Natural Specialized Structured
Params(M)
Average
Cifar100
Caltech101
DTD
Flowers102
Pets
SVHN
Sun397
Camelyon
EuroSAT
Resisc45
Retinopathy
Clevr-Count
Clevr-Dist
DMLab
KITTI-Dist
dSpr-Loc
dSpr-Ori
sNORB-Azim
sNORB-Ele
Conventional Fine-Tuning
FFT-Vim 26 70.1 47.7 89.4 64.2 89.0 87.7 90.6 35.1 84.5 93.9 81.0 74.5 67.5 52.9 47.3 78.9 75.3 53.9 33.3 29.4
FFT-ViT 86 69.9 49.4 89.3 65.5 91.7 89.1 91.4 33.5 85.9 93.6 85.4 74.3 54.7 55.2 48.7 79.7 68.2 49.7 31.5 27.7
LP-Vim 0 55.3 40.9 83.3 57.3 66.3 86.3 38.4 34.6 79.0 87.6 65.0 73.6 36.3 35.1 33.3 64.8 23.0 21.6 15.1 21.7
LP-ViT 0 66.4 50.6 85.6 61.4 79.5 86.5 40.8 38.0 79.7 91.5 71.7 65.5 41.4 34.4 34.1 55.4 18.1 26.4 16.5 24.8
PEFT on ViT
AdaptFormer 0.147 72.4 56.2 89.6 67.2 91.2 91.1 85.9 42.1 85.4 94.6 84.0 74.3 75.8 58.6 48.6 79.6 81.6 53.7 29.6 35.2
LoRA 0.295 72.5 56.4 89.0 66.9 91.2 90.4 86.9 41.5 85.4 95.1 84.1 75.2 75.8 61.7 47.7 80.5 80.4 52.0 29.4 35.7
PEFT on Vim
LoRA 0.054 70.1 57.5 87.7 64.4 86.0 90.0 85.7 39.8 82.2 93.8 79.6 72.5 78.6 56.5 42.0 80.5 71.8 51.0 28.4 32.6
Lily-S 0.074 71.4 58.2 88.5 65.6 87.1 90.7 87.5 40.4 83.3 94.1 79.7 73.8 81.2 57.3 44.1 80.9 79.3 54.1 30.0 33.7
Lily-L 0.196 72.3 57.8 89.4 66.2 87.8 90.5 88.1 40.5 84.1 94.3 81.3 75.1 81.6 57.8 46.5 81.0 82.9 55.2 32.1 34.8
Implementation We assess Lily on the Visual Task Adaptation Benchmark (VTAB-1K Zhai et al. (2019)), a suite of 19 visual tasks spanning diverse domains and semantics, to test its general visual adaptation capability. Tasks are categorized into Natural, Specialized, and Structured, all formulated as classification problems for consistent model evaluation. We conduct two sets of experiments: one focusing on the adaptation effectiveness on Vision Transformer (ViT (Dosovitskiy et al., 2020)) and the other on Vision Mamba (Vim (Zhu et al., 2024)), demonstrating Lily’s architecture-agnostic capabilities. For ViT, we use ViT-B pre-trained on ImageNet-21K (Deng et al., 2009), and for Vim, Vim-s pre-trained on ImageNet-1K. To fairly compare ViT and Vim architectures, we implement LoRA (Hu et al., 2021) and AdaptFormer (Chen et al., 2022) on ViT-B pre-trained on ImageNet-1K. In ViT experiments, we compare Lily with LoRA, AdaptFormer, FourierFT (Gao et al., 2024b), and MoRA (Jiang et al., 2024); in Vim experiments, we focus on contrasting architectures difference, therefore only using LoRA as the baseline. All experiments include FFT and linear probing as baselines. For Vim, we implement two versions: Lily-S (Small) and Lily-L (Large) of Lily, with different hyperparameter settings to either reduce the parameter count (Lily-S) or maximize performance (Lily-L). For Lily on ViT, the reported results are obtained from adapting both the self attention and the MLP module in Transformer. For the performance w.r.t the fine-tuned module, we conduct additional experiments in Appendix D. Detailed experimental settings and datasets information are provide in Appendix B.1.3 and B.2.3.
Results Results are shown in Table 4 and Table 5. For ViT, Lily significantly outperforms all compared PEFT methods with improved parameter efficiency. For Vim, results on ViT generally surpass those on Vim. For instance, LoRA on ViT performs better than LoRA on Vim. We argue that this is due to differences in architecture designs and general model sizes. However, Lily’s strong adaptation performance allows it to match or exceed PEFT methods on ViT and significantly outperform LoRA on Vim (Lily-S and Lily-L surpass LoRA by a significant margin). This demonstrates Lily’s architecture-agnostic capability, highlighting its potential across various model architectures. In general, Lily has achieved great visual adaptation capability with an advantage of being architecture-agnostic and enjoying excellent parameter-efficiency.
Figure 3: Attention maps of Lily and LoRA. The input images for the example here are taken from Caltech101 datasets from VTAB-1K benchmark. It can be observed that features from a certain layer have more similarity to those in other layers in Lily than in LoRA.
Figure 4: Feature difference measured in absolute distance for each element. We compare Lily and LoRA in terms of the difference between features from different layers. In this example image taken from Caltech101, we visualize the feature difference between layers 6 and 1, as well as between layers 6 and 9.
4.5 Understanding Lily
From a Feature Merging Perspective. Apart from being a better gradient projector than LoRA, Lily also enables comprehensive information access across layers. Lily enables access to information or features from all other layers when adapting a target module at a specific layer thanks to the inter-connectivity of the adapters. We aim to understand how Lily achieves this comprehensive information access from the perspective of visual tasks as shown in Fig. 3. We can observe that, in Lily, the distinctness of the attention maps between layers is not as pronounced as in LoRA. This validates Lily’s ability to enable all-level information access, since adaptation at each layer takes into account features from other layers. Additionally, we specifically visualize the actual feature differences between different layers in Fig. 4. We observe that Lily has more points with low feature differences (blue color) than LoRA, indicating that the distinctness of features between layers in Lily is generally lower than in LoRA. This further demonstrates Lily’s ability to enable comprehensive information access. Although we enable all-level information access, what prevents the features from becoming completely identical is the selectivity introduced by Lily, which we specify in the following section.
Figure 5: Impact of attention granularity (i.e., the choice of how many LPs and HPs) on the performance. We choose 12 out of 19 tasks from VTAB-1K for a comprehensive understanding.
Figure 6: Visualization of accumulated assigned weight for HP experts by a router across various layers. Example here uses layer of index 2, 13 and 22 to represent shallow, middle and deep layers.
What’s the Influence of Attention Granularity? The number of experts in the model-wide HP module can be freely set, and the number of LPs can also be flexibly set by sharing across the same level of layers introduced in Appendix A.1. Therefore, we analyze the impact of these choices on performance. We denote the number of LP experts and HP experts as ne_1 and ne_2, respectively. For simplicity, we make them identical in the experiments, denoted as ne. We refer to the number of layers each expert attends to as attention granularity. As the value of ne increases, the attention granularity becomes finer. As shown in Fig. 5, the results from the VTAB-1K benchmark indicate different patterns. For instance, on the DTD dataset, the best performance is achieved when ne is 4, while on sNORB-Azim, performance increases with the increase in ne. Increasing ne leads to more parameters and finer attention granularity. However, finer attention granularity does not necessarily lead to better overall performance. For example, on Resisc45, DTD, Cifar100, sNORB-Ele, dsPr-LoC, Flowers102, and EuroSAT, the negative impact of increasingly finer attention granularity eventually outweighs the benefits of increased parameters, leading to a decrease in overall performance. In other tasks, different patterns may occur because the positive effect of attention granularity on performance is consistently strong, or its negative effect is not enough to offset the benefits of increased parameters, resulting in a generally increasing performance with ne. This phenomenon provides an important insight: for most tasks, simply increasing parameters may not lead to better performance. Instead, only when attention granularity and the number of parameters reach a good tradeoff can we achieve the best performance.
Does Lily Exhibit Selectivity? Lily uses routers to assign varying weights to different HP experts, thereby achieving selective information combination. We illustrate this selectivity in Fig. 6. We use a setup with three HP experts and select three layer levels (1, 13, 22) to calculate the total weight assigned to each expert. The results reveal a clear selectivity: for different layers, the router assigns significantly different weights to different HP experts. For instance, on Cifar100, the middle layer is predominantly dominated by HP 2, whereas the deep layer is primarily dominated by HP 1 and HP 2. In contrast, on Retinopathy, both the middle and deep layers are dominated by HP 3. This selectivity ensures that, even when different layers share information, the inherent differences between layers are still taken into account, making the adaptation more flexible and comprehensive.
5 Conclusion
In this paper, we propose low-rank interconnected adaptation (Lily), a novel framework for efficient fine-tuning via inter-connectivity of adapters. Lily enables each layer to access information from others during adaptation through a hierarchical structure. Additionally, it successfully overcome the low-rank update limitation of LoRA, enabling high-rank update and therefore better adaptation capability. Our approach consistently improves performance across various modalities, model sizes, and architectures, surpassing existing methods with enhanced efficiency. In summary, Lily’s versatility and efficiency make it a promising approach for a wide range of applications.
Appendix
Appendix A More discussion about Lily
A.1 Model Structure and Design Intuition of Lily
Within the overall framework of Lily, we delve into specific implementation details and model design insights. First, we have established the relationship between LPs and HP: LP is confined to specific levels of layers, capturing features that enable the router to selectively assign weights to the HP experts. In contrast, HP is a model-wide module comprising multiple experts, each of which contains information from a particular level of layers. We highlight several key aspects which are not heavily discussed in the methodology section:
A.1.1 Number of LPs
Since LP is limited to specific layers, the simplest approach would be to place an LP at each layer of the module to be adapted (e.g., the query transformation in MHSA). However, this setup may not be necessary: the importance of each layer varies, and many layers have significantly lower importance than others [Zhang et al., 2023]. To achieve greater parameter efficiency, we can set up fewer LPs, with each LP focusing on a level of layers rather than a single layer. For example, an LP can focus on shallow layers (e.g., layers 0, 1, 2, etc.) or deep layers. To enable a single LP to handle multiple layers, we can share an LP across multiple layers. By doing so, we eliminate the redundancy of having an LP at each layer, reduce the number of parameters, and increase efficiency. This is exactly the strategy adopted by most of the experiments.
A.1.2 Number of HP Experts
Regarding HP, the number of experts can be arbitrarily set, enabling more flexible configurations. In our experiments, for the sake of simplicity, we set the number of HP experts equal to the number of LPs, thereby equating the granularity of LP and HP.
A.1.3 Routers Setup
There are also different settings that can be employed for the router. First, we can bind the router to HP, resulting in only one router per model. However, since the number of parameters in the router is relatively small, having only one router per model may not lead to significant selectivity. Therefore, we can also bind the router to LP, configuring a separate router for each LP. Most of our experiments use the latter setup, but in the vision experiments on Vim, we use the single-router and no-lp-sharing setup to verify its effectiveness. From the results, we can see that this setup also performs well. As future work, we can verify the effectiveness of using the latter setup on Vim, which may potentially lead to superior performance.
A.1.4 Hyperparameters
We detail the hyperparameters used in Lily. Specifically, we use lily_r to represent the hidden-dimension of the projectors: LPs and HPs. It serves the same function as r in LoRA. We use lily_s to represent the scaling factor used by Lily. It is mostly searched within the range of {0.01, 0.1, 1.0, 10.0, 100.0}. We use ne_1 to represent the number of LPs used in the model. Since the LPs can be shared as discussed in the previous section, ne_1 does not need to equal the number of layers in the model. We use ne_2 to represent the number of HP experts in the model-wide HP module. In our experiments, we set ne_1 = ne_2 to improve parameter-efficiency and simplicity.
A.1.5 Design Intuition
Lily employs a hierarchical structure to enable updates with higher-ranks than LoRA. However, simply equally connecting all the HPs to the LPs can not achieve the best performance. From the perspective of feature and information utilization across layers, simply aggregating all HPs for an LP ignores the distinctness of the features from current layers. Meanwhile, it reduced the variability of the combinations of gradient projection matrices ($S_{i}$ and $C_{i,j}$ are constants now), making the rank of the weight update higher than that of LoRA (since multiple distinct random matrices are used), but still not high enough for the best performance because of the lack of variability during combination. Therefore, we introduce selectivity into the inter-connectivity as well as discussed below, making the combination of HPs data-dependent so that each $S_{i}$ is unique across the time-steps, enabling updates with even higher ranks. The approach has similarity to Hao et al. [2024], where random matrix is constantly resampled to ensure updates with higher ranks. We further conduct an analysis in Appendix G.
A.2 Efficient Implementation for Weighted Combination
One intuitive implementation of the weighted combination in Lily is Eq. 17, from which we could observe that it perform $N_{e}$ times of matrix multiplication, $N_{e}$ times of scalar multiplication and $N_{e}$ times of matrix addition. Therefore, despite its intuitive nature, the computational burden of this approach is quite formidable.
However, Eq. 18 which is adopted in Lily only utilize $N_{e}$ times of scalar multiplication, $N_{e}$ times of matrix addition and 1 time of matrix multiplication. This saves roughly $N_{e}$ times of matrix multiplication, which can be significant as the size of the model and number of adaptation targets increases. For a $x^{\prime}$ size of $\mathbb{R}^{N\times d}$ and a $P_{H}\in\mathbb{R}^{d\times C}$, the floating-point operation (FLOPs) of these two implementation are:
FLOPs $=\sum_{1}^{N_{e}}(2NdC)+\sum_{1}^{N_{e}}(dC)+\sum_{1}^{N_{e}}(NC)$ (21)
$=N_{e}\times(2NdC+dC+NC)$ (Intuitive)
FLOPs $=2\sum_{i=1}^{N_{e}}(dC)+2NdC$
$=2dC\times(N+N_{e})$ (Lily)
from which we can easily observe that the approach adopted by Lily requires less computation and therefore provides more speed and efficiency during the fine-tuning process. Under the setting of $N=1024,d=16,C=768,N_{e}=4$, the FLOPs of the intuitive approach would be $0.104$ GFLOPs while in Lily it is merely $0.025$ GFLOPs, which could potentially lead to a 4X increase in speed.
A.3 Actual implementation of Lily
Figure 7: Implementation of Lily in VTAB-1K benchmark.
We present the actual implementation of Lily in Fig. 7. For the example here, we choose the implementation from visual adaptation tasks (i.e., VTAB-1K benchmark). For LLM, the implementation is a bit more complicated because of modifications to the huggingface PEFT library [Mangrulkar et al., 2022], but the fundamental adaptation process is the same. Specifically, given an input, we first use the corresponding LP of the current layer to project it to a low-dimensional representation. After that, we use the low-dimensional representation to selectively assign weights for the HP experts. Once we obtain all the weights for the experts, we set out to combine these HP experts accordingly, as discussed in Appendix A.2. After the weighted combination, we use the obtained combined HP to project the low-dimensional representation to high-dimension, therefore acquiring the extra knowledge gained through adaptation.
Appendix B Experimental settings
B.1 Hyper-parameters
A detailed description of the hyper-parameters used in Lily is in Appendix A.1.
B.1.1 Commonsense Reasoning
The hyper-parameters used in commonsense reasoning experiments for MiLoRA, PiSSA are provided in Table 7, 6. Settings for Lily and LoRA using Falcon-Mamba as the backbone are provided in Table. 9 and 8. It can be noticed that Lily achieves the best performance by merely adapting the multi-head self attention module (MHSA) in LLaMA3-8B, while other compared methods adapt all the modules including MLP. Meanwhile, Lily employs the least amount of parameters, showcasing its excellent adaptation at low parameter-budget scenarios.
Table 6: Hyperparameter configuration from the MiLoRA paper.
MiLoRA hyperparameters
Rank r 32
$\alpha$ of LoRA 64
$\alpha$ of PiSSA 32
Dropout 0.05
Optimizer AdamW
LR 3e-4
LR Scheduler Linear
Batch Size 16
Warmup Steps 100
Epochs 3
Placement query, key, value, MLP up, MLP down
Table 7: Hyperparameter configuration from the PiSSA paper.
PiSSA hyperparameters
$\alpha$ Same as rank r
Dropout 0.0
Optimizer AdamW
LR 2e-5
LR Scheduler cosine
Batch Size 128
Warmup Ratio 0.03
Epochs 1
Placement query, key, value, output, gate, MLP up, MLP down
Table 8: Hyperparameter configuration for LoRA using Falcon-Mamba as backbone.
LoRA hyperparameters
Rank r 2
$\alpha$ 16
Dropout 0.05
Optimizer AdamW
LR 3e-4
LR Scheduler Linear
Batch Size 16
Epochs 1
Placement input, delta
Table 9: Best Hyperparameter configuration for Lily using Falcon-Mamba and LLaMA3 as backbones.
Falcon-Mamba LLaMA3
Rank r 40 16
ne_1 4 4
ne_2 4 4
Dropout 0 0
Optimizer AdamW AdamW
LR 3e-4 3e-4
LR Scheduler Linear Linear
Batch Size 16 16
Epochs 1 3
Placement input query, key, value
Table 10: Hyperparameter of Lily on GLUE benchmark.
Hyperparameter STS-B RTE MRPC CoLA SST-2 QNLI MNLI QQP
Optimizer AdamW
LR Schedule Linear
Learning Rate (Lily) 5e-3
Learning Rate (Head) 5e-3
Max Seq. Len 512 512 512 512 512 512 512 512
Lily_s 0.1 0.1 0.1 0.01 0.01 0.01 0 0
ne_1 2 3 2 4 2 2 0 0
ne_2 2 3 2 4 2 2 0 0
Batch Size 64 32 50 64 32 32 0 0
B.1.2 Natural Language Understanding
Specific hyper-parameter settings of Lily on GLUE benchmark are provided in Table 10. We fix the learning rate of both the backbone and the head as 5E-3 and tune the scaling factor lily_s $\in\{0.01,0.1,1.0\}$ instead. For the rank $r$ we fix it to 32 and the seed to 0.
Table 11: Hyperparameter configuration for Lily on VTAB-1K benchmark.
Vision Transformer Vision Mamba
Optimizer AdamW AdamW
Batch Size 64 64
Learning Rate 1E-3 1E-2
Weight Decay 1E-4 1E-3
# Epochs 100 100
LR Decay cosine cosine
B.1.3 Visual Adaptation Benchmark
We provide the hyper-parameter for Lily on VTAB-1K benchmark in Table 11. Specifically, we fix the learning rate at 1E-3 with a weight decay of 1E-4. For ViT, we tune the scaling factor lily_s $\in\{0.01,0.1,1.0,10.0\}$ to maximize the performance, following Jie et al. [2023], Jie and Deng [2023]. For Vim, we fix lily_s to 1.0. Additionally, we search for the hyper-parameters ne_1 and ne_2 within the range {2, 3, 4}, as these numbers can divide the number of layers in the ViT model (12 in ViT-B). For vim, we use the implementation discussed in section A.1, which does not share LPs across layers. Therefore, ne_1 in this setting is fixed to number of layers in Vim (22 in this case), while we search ne_2 in {3,6} and {5,6,17} separately for Lily-S and Lily-L. Note that ne is only set for input projection in Vim. For delta transformation, we only use a single HP expert to reduce the parameter cost. In the experiments of ViT, the rank $r$ is fixed at 16. Meanwhile in Vim’s setting, we tune the ranks $r$ for the delta transformation module and the input projection module separately. We use $4,4$ and $4,8$ separately for Lily-S and Lily-L.
B.2 Datasets
B.2.1 Commonsense Reasoning
Table 12: Details of the datasets used in our commonsense reasoning tasks.
Benchmark Description # Test Questions
ARC-c Multiple-choice science 2376
ARC-e Multiple-choice science 1172
OBQA Multi-step reasoning 500
SIQA Social implications 1954
WinoG Fill-in-a-blank 1267
PIQA Physical commonsense 1830
BoolQ Yes/no questions 3270
HellaS Commonsense NLI 10042
We provide a short description of each datasets used in commonsense reasoning experiments in Table 12.
B.2.2 Natural Language Understanding
We provide detailed information about datasets in the GLUE benchmark in Table 13.
Table 13: Information about datasets in the GLUE benchmark, with STS-B being a regression task and all other tasks falling into the categories of single-sentence or sentence-pair classification.
Corpus Metrics Task # Train # Val # Test # Labels
Single-Sentence Tasks
CoLA Matthews Corr. Acceptability 8.55k 1.04k 1.06k 2
SST-2 Accuracy Sentiment 67.3k 872 1.82k 2
Similarity and Paraphrase Tasks
MRPC Accuracy/F1 Paraphrase 3.67k 408 1.73k 2
STS-B Pearson/Spearman Corr. Sentence similarity 5.75k 1.5k 1.38k 1
QQP Accuracy/F1 Paraphrase 364k 40.4k 391k 2
Inference Tasks
MNLI Accuracy NLI 393k 19.65k 19.65k 3
QNLI Accuracy QA/NLI 105k 5.46k 5.46k 2
RTE Accuracy NLI 2.49k 277 3k 2
B.2.3 Visual Adaptation Benchmark
We provide detailed information about all the tasks from VTAB-1K benchmark in Table 14.
Table 14: Detailed information about the datasets in VTAB-1K benchmark.
Dataset Train Val Test #Classes
VTAB-1k CIFAR100 800/1,000 200 10,000 100
Caltech101 6,084 102
DTD 1,880 47
Oxford-Flowers102 6,149 102
Oxford-Pets 3,669 37
SVHN 26,032 10
Sun397 21,750 397
Patch Camelyon 32,768 2
EuroSAT 5,400 10
Resisc45 6,300 45
Retinopathy 42,670 5
Clevr/count 15,000 8
Clevr/distance 15,000 6
DMLab 22,735 6
KITTI-Dist 711 4
dSprites/location 73,728 16
dSprites/orientation 73,728 16
SmallNORB/azimuth 12,150 18
SmallNORB/elevation 12,150 18
Appendix C Does Sharing LP Results in Inferior Performance?
As mentioned earlier, we adopted a strategy of sharing the LP across most of our experiments, ensuring that the number of LP and HP experts is consistent. This approach offers two benefits: simplicity and enhanced parameter efficiency. By sharing the LP, we eliminate the need to set a separate LP for each layer, thereby reducing the overall parameter count.
Our decision to share the LP is based on the observation of overall redundancy among layers. Specifically, different layers have varying levels of importance [Zhang et al., 2023], and some less important layers do not require a dedicated LP. By not setting a separate LP for these layers, we avoid introducing extra parameter overhead while having a negligible impact on performance. To test that whether sharing LP results in inferior performance, we conduct experiments with no LP sharing on VTAB-1K. The results are shown in Table 15, from which we can observe that the best overall performance ($77.3\%$) is the same as that in the LP-sharing setting. This indicates that even if we employ one LP for each layer, the performance gain is negligible and many of the parameters are actually redundant. However, not sharing LPs results in extra parameter overhead and damages the parameter-efficiency of Lily. Therefore, LP-sharing is a great strategy to eliminate redundancy among LPs and boost the parameter-efficiency of Lily.
Appendix D Where to Apply Lily in Transformers?
Table 15: Performance on VTAB-1K benchmark when applying Lily to various modules in Transformer. The implementation here does not share LP for simplicity (i.e., each layer has one LP).
Natural Specialized Structured
Average
Cifar100
Caltech101
DTD
Flowers102
Pets
SVHN
Sun397
Camelyon
EuroSAT
Resisc45
Retinopathy
Clevr-Count
Clevr-Dist
DMLab
KITTI-Dist
dSpr-Loc
dSpr-Ori
sNORB-Azim
sNORB-Ele
qv 76.9 73.2 92.3 72.2 99.3 91.4 89.0 56.5 87.6 95.2 84.8 75.9 83.7 65.8 52.8 81.2 87.6 52.4 36.3 43.4
mlp 77.0 74.0 92.6 72.2 99.4 91.5 89.0 55.9 88.2 95.5 85.4 76.0 83.3 63.1 53.0 81.4 86.5 53.8 35.6 43.3
qvmlp 77.1 73.9 93.2 72.7 99.4 91.6 89.7 56.5 87.9 95.3 85.0 76.1 84.6 65.2 53.0 82.1 86.7 53.0 36.0 42.8
kvmlp 77.3 74.0 92.3 72.6 99.3 91.5 89.2 56.7 88.2 95.4 85.3 76.0 84.6 64.9 53.4 81.7 87.5 52.9 36.9 45.2
PEFT methods have been predominantly explored on the Transformer architecture, which consists of multi-head self-attention (MHSA) and multi-layer perceptron (MLP) as its core modules. In this section, we analyze the impact of fine-tuned modules on performance using Lily. Specifically, we compare Lily’s performance on the VTAB-1K benchmark under four settings:
  • Applying Lily solely to the query and value transformation module in MHSA (denoted as "qv").
  • Applying Lily solely to the MLP module (denoted as "mlp").
  • Applying Lily to both the query and value transformation module in MHSA and the MLP module (denoted as "qvmlp").
  • Applying Lily to both the key and value transformation module in MHSA and the MLP module (denoted as "kvmlp").
To ensure a fair comparison, we tune the hyperparameters to maintain a similar parameter count across all settings. Additionally, to further investigate whether sharing the low-rank projection (LP) affects performance, we do not share the LP in this experiment. The results are presented in Table 15. We observe that the "kvmlp" setting achieves the best performance, with an average accuracy of $77.3\%$. In contrast, adapting only the MHSA module ("qv") yields the worst performance. Furthermore, we note that adapting both the MHSA and MLP modules (qvmlp and kvmlp) generally leads to superior results compared to adapting only one specific module (qv and mlp). This suggests that both MLP and MHSA play crucial roles in the overall model performance, and adapting both is essential for effective adaptation.
Notably, even when applying Lily solely to the MHSA module, which results in the worst performance among the four settings ($76.9\%$), it still outperforms LoRA by a significant margin ($0.5\%$). This underscores the efficiency of Lily, as it uses fewer parameters than LoRA even without LP sharing.
Appendix E Where to Apply Lily in Mamba?
Nearly all previous PEFT method studies have been centered around Transformers, while Mamba is a relatively new architecture, so there has been little research on PEFT methods on Mamba. In this section, we briefly analyze the pros and cons of adapting Mamba’s modules. In brief, a Mamba block consists of regular linear projection layers and a core component SSM module [Gu and Dao, 2023] [Zhu et al., 2024]. Specifically in SSM module, Mamba utilize parameters ($\Delta$, $A$, $B$, $C$) to transform an input sequence $x(t)$ to an output sequence $y(t)$ using a hidden state $h(t)$. The discretization process converts $A$ and $B$ into $\bar{A}$ and $\bar{B}$, respectively, using the time step size parameter $\Delta$. Structured state space models, inspired by continuous systems, can be computed similarly to RNNs or in the form of global convolution due to their linear time invariance (LTI) property. Mamba introduces a selective property to structured state space model, tying parameters to the current input, which breaks the LTI property and hinders parallel training. To address this, Mamba employs a hardware-aware algorithm, enabling its SSM module to possess the selective property and perform parallel training. To be specific, the discretization process can be expressed as:
$\bar{A}$ $=exp(\Delta A)$ (22)
$\bar{B}$ $=(\Delta A)^{-1}(exp(\Delta A)-I)\cdot\Delta B$
After that, the calculation in Mamba can be expressed as:
$h_{t}$ $=\bar{A}h_{t-1}+\bar{B}x_{t}$ (23)
$y_{t}$ $=Ch_{t}$
where $h_{t}$ is the hidden state at time $t$ and $x_{t}$ is the corresponding input token. Delta projection is a module in SSM that’s learnable and tasked with transforming the parameter $\Delta$. Since adapting the delta projection alone can indirectly adapt the entire SSM module (i.e., $\bar{A}$ and $\bar{B}$ are all determined by $\Delta$), it is the most critical component of the SSM module.
We investigate the performance of two adaptation strategies: adapting only the input linear projection layer (denoted as "in") and adapting both the input linear projection layer and SSM (denoted as "$\Delta$ + in" since we only adapt delta projection in SSM module). Our results, as shown in Table 1, indicate that applying Lily solely to the input projection yields better performance than applying it to both the input and delta projection modules. This suggests that when adapting Mamba-based models under the paradigm of low-rank adaptation, it is optimal to adapt only the input projection module outside the SSM module. These findings highlight the need for further research into the impact of fine-tuned modules in Mamba on overall performance. Additionally, developing PEFT methods specifically tailored to Mamba-based models, whether for vision or language foundation models, is also a promising direction for future work.
Appendix F Performance with Different Learning Rates
Since we only tuned the learning rate in the commonsense reasoning experiment, we provide the performance of commonsense reasoning under different learning rates in Table 16.
Table 16: Commonsense reasoning results of Lily under various leanring rates.
Model Lr BoolQ PIQA SIQA HellaSwag WinoGrande ARC-e ARC-c OBQA Avg.
LLaMA3-8B 1e-3 70.7 84.6 77.6 87.8 77.3 88.5 74.1 80.8 80.2
5e-4 71.8 86.5 77.9 82.8 83.1 88.6 76.8 81.4 81.1
3e-4 72.9 85.6 77.8 92.7 83.3 89.7 77.6 82.8 82.8
Appendix G Does Selectivity Help?
Figure 8: Implementation of Lily with no selectivity.
Lily introduced selective weight combination to selectively incorporate information from other layers. To verify the effectiveness of this selectivity, we remove the router from Lily and evaluate the impact. The modified algorithm without the router is presented in Fig. 8. We conduct experiments on commonsense reasoning to investigate the effect of removing selectivity from Lily.
As shown in Table 17, removing selectivity from Lily results in generally poorer performance compared to vanilla Lily. This is likely because the lack of selectivity causes Lily to simply aggregate all the HP expert, leading to inferior performance. This validates the design choice of using routers in Lily to selectively allocate weights to HP experts, rather than simply summing them.
Table 17: Commonsense reasoning results of Lily without selectivity. We provide results using two learning rates.
Model Lr BoolQ PIQA SIQA HellaSwag WinoGrande ARC-e ARC-c OBQA Avg.
LLaMA3-8B 3e-4 64.0 82.6 78.5 77.0 79.6 88.4 74.5 82.0 78.3
5e-4 71.3 85.5 78.1 84.3 79.6 86.4 76.1 79.0 79.8
Appendix H How to Allocate Parameters?
Since Lily alters the traditional LoRA’s layer-bound setup, increasing the parameters of Lily can be achieved through two approaches: 1) increasing ne, i.e., increasing the number of LP and HP experts, and 2) increasing the rank, i.e., increasing the parameter size of each individual LP or HP expert. In this section, we investigate which factor has the greatest impact on performance. We conduct experiments on the commonsense reasoning task. Specifically, we maintain the same parameter count and learning rate, and achieve the same parameter count by setting different ranks and adjusting the corresponding ne (e.g., r=16, ne=4 versus r=8, ne=8). The results are shown in Fig. 9, from which we observe that more LP and HP experts with smaller rank (i.e., bigger ne and smaller rank) generally performs worse. We argue that this is because, although increasing the attention granularity allows for finer details, the resulting performance gain is not as significant as the gain obtained by increasing the rank, i.e., increasing the model’s capacity to learn more information. This gives us an insight that, in Lily, increasing ne to increase the parameters is less effective than directly increasing the rank in terms of potential performance gain.
Figure 9: Results on commonsense reasoning tasks when applying different settings of rank. The hyperparameter ne is specifically tuned to maintain the same amount of parameter count for a fair comparison.
Appendix I More on Subject-driven Generation
We provide more results on subject-driven generation in Fig. 10 and Fig. 11.
Figure 10: More subject-driven generation results for unreported subjects.
Figure 11: More subject-driven generation results for subjects that are reported in the experiment section.
Appendix J More on Attention Maps of Lily and LoRA
We provide more visualization results of the attention map from both LoRA and Lily on Caltech101 dataset from VTAB-1K benchmark in Fig. 12.
Figure 12: More results of attention maps from LoRA and Lily. All images are taken from Caltech101 dataset.

### v3

@@
 Abstract
-Low-rank adaptation (LoRA) is a powerful parameter-efficient fine-tuning method that utilizes low-rank projectors $A$ and $B$ to learn weight updates $\Delta W$ for adaptation targets $W$.
-Previous research has shown that LoRA is essentially a gradient compressor, performing random projections on the gradient using a fixed projection matrix $A_{0}$.
-However, this setup restricts the overall weight update to be low-rank, which limits the adaptation performance.
-In this paper, we propose low-rank interconnected adaptation across layers (Lily).
-Specifically, we employ a hierarchical framework where low-dimensional projectors (LPs) retained for downward projection at a particular level, while globally-shared high-dimensional projector (HP) experts perform upward projection across all levels of layers.
-Lily uniquely connects each LP to all HP experts, therefore the gradient projections are no longer dominated by fixed projection matrices, but rather by selective combinations of all the projectors, thereby breaking the low-rank constraint of LoRA.
-Furthermore, Lily’s cross-layer connections facilitate the capture of intricate information and dependencies across different layers, thereby enhancing the model’s representational capabilities.
-Experiments across various modalities, architectures, and model sizes underscore Lily’s great performance and efficiency.
-Code is available on github https://github.com/yibozhong/lily.
+[…]
+Low-Rank Interconnected Adaptation across Layers
+Yibo Zhong1, Jinman Zhao2, Yao Zhou1thanks: Corresponding authorthanks: This work was supported by National Natural Science Foundation of China (Grant 62376172), 1Sichuan University, 2University of Toronto, zhongyibo@stu.scu.edu.cnyaozhou@scu.edu.cn
 1 Introduction
-For foundation models like Transformers (Vaswani et al., 2017b), fine-tuning on downstream tasks is a typical usage, but full fine-tuning (FFT) of large models like large language models (LLMs) incurs huge computational and storage costs and risks forgetting previously learned knowledge (Biderman et al., 2024).
-Linear probing, which fine-tunes only the final modules like classification heads, addresses these issues but leads to significant performance degradation since it doesn’t update weights from the backbone.
-To tackle these challenges, parameter-efficient fine-tuning (PEFT) has received significant attention.
-In PEFT, a model’s backbone weights are frozen, and lightweight trainable modules are introduced to efficiently learn task-specific knowledge.
-Among all PEFT methods, Low-rank Adaptation (LoRA (Hu et al., 2021)) is one of the most widely applied techniques, especially in LLMs.
-LoRA introduces a pair of low-rank projection matrices for each adaptation target, consisting of a downward adapter $A$ and an upward adapter $B$, to approximate $\Delta W$ in FFT.
-Due to its low-rank nature, LoRA offers significant computational and storage savings, effectively alleviating the burdens of FFT while significantly outperforming linear probing by learning the weight updates for backbone weight.
-However, LoRA and many subsequent improvements to the method (Miles et al., 2024), (Zhang et al., 2023), (Zhong et al., 2024) have a limitation: the overall learned weight updates $\Delta W$ are restricted to be low-rank, which limits the model performance during adaptation.
-Previous work (Hao et al., 2024) has shown that LoRA adapters are secretly gradient compressors.
-In particular, LoRA updates to the weight can be approximated by performing random projection on the gradient, using a fixed random projection matrix $A_{0}$, where $A_{0}$ is the initial weight for $A$.
-Since this limitation arises because the projection matrix is always fixed in LoRA, using multiple distinct matrices and their variations to perform the gradient projection could avoid restricting the weight updates to a low-rank subspace.
-This prompts a question: How can we enable high-rank updates by using various projection matrices for an adaptation target?
 Figure 1: Dynamics of LoRA and Lily.
-$N_{l}$ is used to denote the number of layers in the model while $N_{e}$ denoting the number of high-dimensional projector experts from Lily.
-R is representing the router from Lily.
-In this paper, we propose Low-rank interconnected adaptation across layers (Lily), a novel framework for more expressive and performative PEFT.
-Specifically, we decouple the downward low-dimensional projector (LP) and its corresponding upward high-dimensional projectors (HP), making them not tightly-bonded.
-Each LP is connected to all the HPs, and vice versa as shown in Fig.
+In this 6-layer example with a fixed overall parameter budget, LoRA allocates the same parameter budget to each layer, resulting in small rank updates for the weights.
+Lily overcomes this by employing a small number of shared adapters with a much larger rank, achieving higher-rank updates while using the same or even a smaller parameter budget.
+Considering the different characteristics, and to make the adaptation more dynamic, the adapters are mixed according to a data-dependent router, represented by $R$.
+Fine-tuning foundation models like Transformers (Vaswani et al., 2017) on downstream tasks is common but costly, especially for large models like LLMs, which incur high computational and storage demands and risk catastrophic forgetting (Biderman et al., 2024).
+Linear probing alleviates these issues by fine-tuning only the final modules, but suffers from performance loss due to frozen backbone weights.
+To address this, parameter-efficient fine-tuning (PEFT) freezes the backbone and introduces lightweight modules for task-specific learning.
+Among PEFT methods, Low-rank Adaptation (LoRA (Hu et al., 2021)) is widely used, particularly for LLMs.
+LoRA introduces low-rank projection matrices, $A$ and $B$, to approximate weight updates $\Delta W$, achieving significant savings in computation and storage while outperforming linear probing by updating the backbone weights.
+However, LoRA and its subsequent improvements (Miles et al., 2024; Zhang et al., 2023; Zhong et al., 2024) face a limitation: the learned weight updates $\Delta W$ are constrained to be low-rank, limiting model performance.
+A key issue is that LoRA allocates the same parameter budget to each layer, regardless of their importance (Fig.
+1).
+As a result, the rank of each adapter is constrained by the fixed budget, raising the critical question: Can we enable more dynamic, expressive adaptation with high-rank weight updates under the same parameter budget?
+In this paper, we propose Low-rank interconnected adaptation across layers (Lily), a novel framework for more expressive and efficient PEFT.
+Specifically, we decouple $A$ and its corresponding upward $B$, eliminating their tight coupling.
+Each $A$ is connected to all $B$s, and vice versa, as illustrated in Fig.
 1.
-This results in a hierarchical structure where LPs are still retained at a particular level to perform downward projection, while all HPs are now globally shared by all the LPs, performing upward projection.
-Inspired by self-attention Vaswani et al.
-(2017a), which calculates the relationship between a token and all tokens and obtains attention scores indicating the strength of their relationship, we selectively connect an LP with the HPs based on layer features.
-The LP extracts features from the current layer, and based on the extracted features, a data-dependent and selective combination of HPs is performed.
-This is realized by utilizing a router (Shazeer et al., 2017) that outputs a unique weight distribution for HP experts, depending on the current input feature, thereby exhibiting selectivity.
-The gradient projection now uses multiple random projection matrices from all the layers and their selective combinations rather than a fixed matrix, thus effectively avoiding the constraint of low-rank updates from LoRA.
-Furthermore, Lily enables a more comprehensive information access by allowing adapters at each layer to access information from other layers, promoting an interconnected and dynamic learning process, where the adapters can collaborate, share learned knowledge and model dependencies across layers.
-Overall, our contributions include:
-  • We propose Lily, a novel PEFT framework that incorporates cross-layer connections of the projection matrices, thereby breaking the restriction of low-rank weight updates in LoRA.
-  • Lily utilizes routers to selectively connect an LP with multiple HP experts, enabling comprehensive information access and therefore expressive adaptation.
-  • Extensive experiments are conducted across various modalities, architectures, and model sizes, highlighting Lily’s great performance and efficiency in diverse scenarios.
+This creates a hierarchical structure where locally-shared $A$s perform downward projections at specific layers, while globally-shared $B$s perform upward projections across all layers.
+To enhance dynamism, we selectively connect each $A$ with $B$s based on layer features.
+$A$ extracts features from the current layer, and a selective mixture of $B$s is performed based on these features, enabled by routers (Shazeer et al., 2017) that generate data-dependent weight distributions for $B$ experts.
+The interconnected structure makes the adaptation process more dynamic and flexible, with rich interactions between adapters.
+By reducing the number of adapters and increasing their rank, Lily achieves higher-rank weight updates than LoRA while using the same or fewer parameters.
+Additionally, Lily enables comprehensive information access and learning by allowing adapters at each layer to collaborate, share knowledge, and model dependencies across layers.
+Our key contributions include:
+  • We propose Lily, a novel PEFT framework that introduces interconnected adapters, effectively overcoming the limitations of low-rank weight updates in LoRA under the same parameter constraints.
+  • Lily utilizes routers to dynamically select and connect an adapter $A$ with multiple adapter $B$ experts, enabling richer information flow and more expressive adaptation dynamics.
+  • Extensive experiments are conducted across diverse modalities, architectures, and model scales, demonstrating Lily’s superior performance and efficiency in a wide range of scenarios.
 2 Related Work
-Parameter Efficient Fine-Tuning.
-Typical usage of foundation models includes pre-training on large datasets and fine-tuning on various downstream tasks.
-Parameter-efficient fine-tuning (PEFT) thus emerges as a promising field, aiming to fine-tuning the model efficiently with minimal parameters while maintaining performance and preserving previously learned knowledge, addressing drawbacks posed by conventional fine-tuning techniques like full fine-tuning or linear probing.
-Current PEFT research can be mainly categorized into two types: 1) adapter-based methods (Hu et al., 2021), (Chen et al., 2022), (Pfeiffer et al., 2020b), (Jie and Deng, 2023) (Houlsby et al., 2019b) and 2) prompt-based methods (Tu et al., 2023b) (Tu et al., 2023a).
-Adapter-based methods introduce lightweight adapters into the Multi-Head Self-Attention (MHSA) or the Feed-Forward Network (FFN) blocks within the Transformer architecture.
-On the other hand, prompt-based methods append trainable tokens as prompts to the input sequence fed to certain parts of the model.
-Among these various PEFT techniques, low-rank adaptation (LoRA (Hu et al., 2021)) stands out as one of the most well-known methods.
-LoRA introduces a pair of projection matrices $A$ and $B$ per adaptation target $W$.
-The low-dimension projector (LP) $A$ projects input $x$ to low-dimension space, and the high-dimension projector (HP) $B$ restores it to its original dimension.
-Multiplying these projection matrices approximates the weight update $\Delta W$ in FFT.
-Recent work (Hao et al., 2024) has shown that LoRA adapters are essentially performing random projection to the gradient using a fixed matrix.
-This restricts the learned weight update to low-rank subspace and thus imitating the model performance.
-Meanwhile, $A$ and $B$ are tightly coupled, therefore the adaptation process only has information access from current layer, without an understanding of information from other layers, which could be beneficial to modeling dependencies across various layers.
-Mixture of Experts.
-Mixture of Experts (MoE) is an active research area that has garnered significant attention, especially in the field of large language models (LLMs).
-Conditional computation, where different parts of the network are activated on a per-example basis, has been proposed to enhance model capability without increasing computation (Davis and Arel, 2013) (Bengio et al., 2013) (Eigen et al., 2013) (Almahairi et al., 2016).
+Parameter Efficient Fine-Tuning Foundation models are typically pre-trained on large datasets and fine-tuned on downstream tasks.
+Parameter-efficient fine-tuning (PEFT) seeks to fine-tune models efficiently with minimal parameters while maintaining performance and preserving learned knowledge.
+It effectively addresses limitations of conventional fine-tuning techniques, like full fine-tuning or linear probing.
+Current PEFT approaches can be divided into two categories: 1) adapter-based methods (Hu et al., 2021; Chen et al., 2022; Pfeiffer et al., 2020a; Jie and Deng, 2023; Houlsby et al., 2019a; Zhong and Zhou, 2025) and 2) prompt-based methods (Tu et al., 2023a, b).
+Adapter-based methods insert lightweight adapters into the Multi-Head Self-Attention (MHSA) or Feed-Forward Network (FFN) blocks of the Transformer architecture, while prompt-based methods add trainable tokens to the input sequence.
+Among these, low-rank adaptation (LoRA (Hu et al., 2021)) is a well-known technique.
+It introduces projection matrices $A$ and $B$ for each adaptation target $W$, where $A$ projects input $x$ to a low-dimensional space and $B$ restores it to the original dimension.
+The product of these matrices approximates the weight update $\Delta W$ in full fine-tuning (FFT).
+However, this limits the update to a low-rank subspace, which may affect performance.
+Additionally, $A$ and $B$ are tightly coupled, restricting the adaptation process to information from the current layer, which may hinder the modeling of dependencies across layers.
+Mixture of Experts Mixture of Experts (MoE) is an active research area that has received significant attention, especially in the field of large language models (LLMs).
+Conditional computation, which activates different parts of the network on a per-example basis, has been proposed to enhance model capability without increasing computation (Davis and Arel, 2013; Bengio et al., 2013; Eigen et al., 2013; Almahairi et al., 2016) .
 The sparsely-gated MoE layer is introduced to implement this idea, consisting of numerous sub-networks (Shazeer et al., 2017).
-A trainable gating network, known as a "router", determines the combination of experts for each example.
-There are already PEFT methods like MoLORA (Zadouri et al., 2023) and MOLA (Gao et al., 2024a) which apply the MoE design concept to PEFT.
-However, these methods simply treat the adapters combined in LoRA as a single expert.
-A concurrent research Wu et al.
-(2024), utilizes LP and HP sub-spaces as the experts but fails to overcome the limitation discussed in previous section.
+A trainable gating network (router) determines the combination of experts for each example.
+There are already PEFT methods like MoLORA (Zadouri et al., 2023) and MOLA (Gao et al., 2024a) that apply the MoE design concept to PEFT.
+However, these methods simply treat the adapters $A$ and $B$ combined in LoRA as a single expert.
+Concurrent research Wu et al.
+(2024) utilizes $A$ and $B$ sub-spaces as the experts but fails to overcome the limitation discussed in the previous section.
+Another concurrent work, HydraLoRA Tian et al.
+(2024), explores an asymmetric design for LoRA.
+Unlike our work, we consider the interconnection across layers and deploy a model-wide asymmetric design to enable cross-layer connections.
+This enables the use of adapters of higher rank than a typical LoRA setup while using the same or fewer overall parameters.
 3 Methodology
-3.1 Preliminaries
-LoRA is a Gradient Projector. Adaptation process of low-rank adaptation (LoRA) can be expressed as follows when setting the rank to $d$, a small hidden-dimension:
-$W=W_{0}+AB$ (1)
-where $W,W_{0}\in\mathbb{R}^{C_{in}\times C_{out}}$ with $C_{in}$ being the input dimension and $C_{out}$ being the output dimension.
-$W_{0}$ is the frozen weight of a target module and $W$ is the actual value of the target during adaptation runtime.
-$A\in\mathbb{R}^{C_{in}\times d}$ and $B\in\mathbb{R}^{d\times C_{out}}$ are low-dimension projectors (LP) and high-dimension projectors (HP), with their product multiplied by a scalar value $s$ (omitted in the equation) to approximate weight update $\Delta W\in\mathbb{R}^{C_{in}\times C_{out}}$.
-We denote the information from $W_{0}$ as the basic knowledge that the pre-trained model already possesses, while $\Delta W=AB$ represents the learned extra knowledge specific to the current downstream task.
-We denote the initial value of $A$ and $B$ as $A_{0}$ and $B_{0}$, and their update as $\Delta A$ and $\Delta B$.
-$B_{0}$ is initialized as 0.
-The approximation is then described as:
-$W$ $=W_{0}+(A_{0}+\Delta A)(B_{0}+\Delta B)$ (2)
-$=W_{0}+A_{0}\Delta B+\Delta A\Delta B$ (3)
-$\approx W_{0}+A_{0}\Delta B$ (4)
-It is proved in Hao et al. (2024) that it is the term $A_{0}\Delta B$ that contributes the most to the weight updates. The forward pass given an input in LoRA is:
-$y=xW=xW_{0}+xAB$ (5)
-where $x\in\mathbb{R}^{C_{in}}$ and $y\in\mathbb{R}^{C_{out}}$ for simplicity. Therefore, the gradient w.r.t. $B$ is:
-$\frac{\partial{\mathcal{L}}}{\partial B}$ $=A^{T}x^{T}\frac{\partial{\mathcal{L}}}{\partial y}$ (6)
-$=A^{T}(\nabla_{W}{\mathcal{L}})$ (7)
-where $x^{T}\frac{\partial{\mathcal{L}}}{\partial y}\in\mathbb{R}^{C_{in}\times C_{out}}$ is the gradient for the weight matrix, which we denote as $(\nabla_{W}{\mathcal{L}})$.
-Using the approximation in Eq.
-2, we can express the adaptation when only updating $B$ using gradient descent:
-$W=W_{0}-\eta\sum_{t=0}^{T}\big{[}A_{0}A_{0}^{T}(\nabla_{W}{\mathcal{L}}_{t})\big{]}$ (8)
-where $T$ is the current time-step and $\eta$ is the adopted learning rate.
-The projection matrix here is $A_{0}$ because $A$ is not updated in the approximation, since its change contribute little to the overall weight updates.
-We can observe that LoRA is essentially performing random projection to the gradient using a fixed random projection matrix $A_{0}$.
-However, this setup limit the weight update $\Delta W$ to be low-rank, which also limit the adaptation performance.
-Lily is a Better Gradient Projector. Consider simply connecting each $A$ with all the $B$s with weights, the forward pass is then:
-$W$ $=W_{0}+\sum_{i=1}^{N_{e}}\Big{[}S_{i}\big{[}(A_{0}+\Delta A)(B_{0}^{i}+\Delta B^{i})\big{]}\Big{]}$ (9)
-$=W_{0}+\sum_{i=1}^{N_{e}}\big{[}S_{i}(A_{0}\Delta B^{i}+\Delta A\Delta B^{i})\big{]}$ (10)
-$\approx W_{0}+A_{0}\sum_{i=1}^{N_{e}}S_{i}\Delta B^{i}$ (11)
-where $N_{e}$ is used to denote the number of $B$s in the model and $S$ is a set of weight scores of current $A$ for all the $B$s, each $S_{i}$ is a scalar value representing the weight for corresponding $B^{i}$.
-Thanks to the inter-connectivity, each $B$ is also connected to all the $A$s. The gradient w.r.t an $B$ is now:
-$\frac{\partial{\mathcal{L}}}{\partial B}$ $=\sum_{i=1}^{N_{l}}C_{i}{A^{i}}^{T}{x^{i}}^{T}\frac{\partial{\mathcal{L}}}{\partial y^{i}}$ (12)
-$=\sum_{i=1}^{N_{l}}C_{i}{A^{i}}^{T}(\nabla_{W_{i}}{\mathcal{L}})$ (13)
-where $N_{l}$ is the number of $A$s in the model and we denote gradient for each corresponding weight matrix as $\nabla_{W_{i}}{\mathcal{L}}$.
-$C$ is a set of weight scores where each $C_{i}$ is a weight score from certain $A_{i}$ for current $B$.
-The forward pass when only updating $B$ can then be expressed as:
-$W=W_{0}-\eta\sum_{t=0}^{T}\Big{[}\sum_{1}^{N_{e}}S_{i}\big{[}\sum_{j=1}^{N_{l}}{C_{i,j}A_{0}A_{0}^{j}}^{T}(\nabla_{W_{j}}{\mathcal{L}}_{t})\big{]}\Big{]}$ (14)
-where $C_{i,j}$ is used to denote weight scores from various $A^{j},j\in\{1,2,...,N_{l}\}$ to $B_{i},i\in\{1,2,...,N_{e}\}$.
-It can be observed that with this setup, the projection to the gradient no longer use a fixed $A_{0}$, but rather utilizing all the projection matrix $A_{0}^{i},i\in\{1,2,...,N_{l}\}$ and their overall complex combinations.
-Note that all the ${A_{0}}^{i}$s are distinct random projection matrices and each weight score ($S_{i}$ and $C_{i,j}$) is unique, data-dependent and thus varies across time-steps.
-Therefore, the weight updates are not restricted to be low-rank, but rather high-rank.
-This setup is precisely what Lily adopts.
-To differentiate from LoRA, we use different notations for the projectors.
-Specifically, we decouple HPs from their original layers, allowing them to form a model-wide shared module of HP experts ($B$ in LoRA), denoted as $P_{H}^{i}$, where $i\in\{1,2,...,N_{e}\}$ and $N_{e}$ is the number of HP experts.
-Meanwhile, we fix LPs ($A$ in LoRA) $P_{L}^{i},i\in\{1,2,...,N_{l}\}$ in each layer for downward projection, where $N_{l}$ is the number of LPs.
-This hierarchical structure makes the projection matrices all interconnected across layers, enabling the modeling of complex dependencies and high-rank updates to the model weights.
-Detailed model structure and design intuitions are discussed in Appendix A.
-3.2 Dynamics of Lily
-Downward Projection and Selective Weight Allocation.
+3.1 Downward Projection and Selective Weight Allocation
 The process is illustrated in the right half of Fig.
 1.
-Initially, we use an LP to project the input $x\in\mathbb{R}^{N\times C_{in}}$ into its low-dimensional representation $x^{\prime}\in\mathbb{R}^{N\times d}$ where $N$ is the sequence length:
-$x^{\prime}=xP_{L}$ (15)
-Inspired by the Mixture of Experts (MoE) paradigm, we employ a router $R\in\mathbb{R}^{N_{e}\times d}$ to selectively assign weights to all HP experts based on their relationship to the current layer’s features ($x^{\prime}$).
-The weight set $S$ is obtained as:
-$S=softmax(\sum_{i=1}^{N}(x^{\prime}R^{T})_{i})$ (16)
-The router selectively combines experts based on the current layer’s features, enabling smart information integration.
-For shallower inputs, the router increases attention for experts specializing in shallow-layer knowledge, while deeper inputs favor experts learning deep-layer knowledge.
-Weighted Combination of Experts and Upward Projection.
-Once we obtain the low-dimensional input $x^{\prime}$ , we combine information from all layers using the model-wide shared global HP module.
-One intuitive approach is to feed $x^{\prime}$ into each HP expert and combine their outputs to obtain the extra knowledge $x_{\Delta}\in\mathbb{R}^{N\times C_{out}}$:
-$x_{\Delta}=\sum_{i=1}^{N_{e}}S_{i}\cdot(x^{\prime}P_{H}^{i})$ (17)
-where $S\in\mathbb{R}^{N_{e}}$ is the set of weight scores for HP experts, obtained through selective weight allocation.
-To address efficiency concerns discussed in Appendix A.2, we propose an alternative implementation that is mathematically equivalent and significantly reduces computational burden, described as:
-$x_{\Delta}=x^{\prime}(\sum_{i=1}^{N_{e}}S_{i}\cdot P_{H}^{i})$ (18)
+Initially, we use an $A$ to project the input $x\in\mathbb{R}^{N\times C_{\text{in}}}$ into its low-dimensional representation $x^{\prime}\in\mathbb{R}^{N\times d}$, where $N$ is the sequence length:
+$x^{\prime}=xA$ (1)
+To enable more parameter efficiency, the number of $A$s can be set to less than the number of layers in the model by sharing the same $A$ across neighboring layers, as illustrated in Fig.
+1 and discussed in A.
+Inspired by the Mixture of Experts (MoE) paradigm, we employ a router $R\in\mathbb{R}^{N_{e}\times d}$ to selectively assign weights to all $B$ experts based on their relationship to the current layer’s features ($x^{\prime}$), where $N_{e}$ represents the number of $B$ experts.
+A weight set $S\in\mathbb{R}^{N_{e}}$ is obtained as:
+$S=\text{softmax}\left(\sum_{i=1}^{N}(x^{\prime}R^{T})_{i}\right)$ (2)
+The router selectively mixes the experts based on this data-dependent weight distribution, enabling information integration and expressive adaptation.
+3.2 Weighted Mixture of Experts and Upward Projection
+Once we obtain the low-dimensional input $x^{\prime}$, we combine information from all layers using the model-wide shared $B$ experts.
+One intuitive approach is to feed $x^{\prime}$ into each $B$ expert and combine their outputs to obtain the additional knowledge $x_{\Delta}\in\mathbb{R}^{N\times C_{\text{out}}}$.
+However, to address efficiency concerns discussed in Appendix A.2, we propose an alternative implementation that is mathematically equivalent but significantly reduces the computational burden, described as follows:
+$x_{\Delta}=x^{\prime}\left(\sum_{i=1}^{N_{e}}S_{i}\cdot B^{i}\right)$ (3)
+where $S$ is the set of weight scores for the $B$ experts, obtained through selective weight allocation.
 Since each $S_{i}$ is a scalar value, the calculation in Eq.
-18 is mathematically equivalent to that in Eq.
-17, but with significantly improved efficiency.
-Therefore, the whole computation flow, with input $x\in\mathbb{R}^{N\times C_{in}}$ and output $y\in\mathbb{R}^{N\times C_{out}}$, for an adaptation target module is:
+3 is mathematically equivalent to the intuitive method but with significantly improved efficiency.
+Therefore, the complete computation flow, with input $x\in\mathbb{R}^{N\times C_{\text{in}}}$ and output $y\in\mathbb{R}^{N\times C_{\text{out}}}$, for an adaptation target module is:
 $y$ $=xW_{0}+s\cdot x_{\Delta}$ (4)
-$=xW_{0}+s\cdot xP_{L}(\sum_{i=1}^{N_{e}}(softmax(\sum_{j=1}^{N}(xP_{L}R^{T})_{j}))_{i}\cdot P_{H}^{i})$ (20)
-By selectively allocating weights and combining HP experts, Lily enables access to all levels of information during adaptation.
-Each layer’s target adaptation modules could consider the status and knowledge from all other layers, resulting in a more expressive and comprehensive adaptation.
-Meanwhile, thanks to its inter-connectivity and selectivity, Lily break the low-rank update constraint of LoRA and enable high-rank updates, as discussed in preliminaries.
+where $s$ is a scaling factor.
+By selectively allocating weights and mixing $B$ experts, Lily enables access to all levels of information during adaptation.
+Each layer’s target adaptation modules can consider the status and knowledge from all other layers, resulting in a more expressive and comprehensive adaptation.
+Meanwhile, thanks to its interconnectivity, Lily can break the low-rank update constraint of LoRA by simply employing a smaller number of adapters with higher ranks.
 4 Experiments
-We validate the effectiveness of Lily across different domains, model sizes (from ViT to LLM), and architectures (Transformers, Mamba), demonstrating its general strong adaptation capability.
-Concurrently, we conduct a comprehensive analysis of Lily’s intrinsic mechanisms, providing a thorough understanding of Lily.
+We validate the effectiveness of Lily across different domains, model sizes (from ViT to LLM), and architectures (Transformers, Mamba), demonstrating its generally strong adaptation capability.
+Concurrently, we conduct a comprehensive analysis of Lily’s intrinsic mechanisms, providing a thorough understanding of how it works.
+All ranks for Lily are selected from 8, 16, 32, ensuring that the total parameter count does not exceed that of the baselines.
 All experiments are conducted on a single RTX 4090 GPU.
-Additionally, multiple analysis are provided in Appendix C, D, E, F, G, H, I and J.
+Additionally, multiple analyses are provided in Appendix C, D, E, F, G, H, I, and J.
 4.1 Common Sense Reasoning
-Table 1: Commonsense reasoning results for Falcon-Mamba-7B across eight tasks. Bold represents the highest performance for each dataset utilizing PEFT methods.
-Model PEFT BoolQ PIQA SIQA HellaSwag WinoGrande ARC-e ARC-c OBQA Avg.
-ChatGPT - 73.1 85.4 68.5 78.5 66.1 89.8 79.9 74.8 77.0
-Falcon-Mamba-7B LoRA 6.5 30.5 40.6 14.9 56.4 42.2 31.8 38.4 32.7
-Lily (Δ + $\Delta\text{ + in}$) 44.9 66.8 65.0 10.5 57.1 78.7 64.6 68.2 57.0
-Lily (in) 60.2 61.0 67.3 12.9 61.5 80.0 67.5 65.8 59.5
+Table 1: Commonsense reasoning results for Falcon-Mamba-7B across eight tasks.
+Bold represents the highest performance for each dataset utilizing PEFT methods.
+“$\Delta$” and “in” refer to adaptations of Mamba’s delta_proj and in_proj parameters, respectively.
+Model PEFT Params BoolQ PIQA SIQA HellaSwag WinoGrande ARC-e ARC-c OBQA Avg.
+ChatGPT - - 73.1 85.4 68.5 78.5 66.1 89.8 79.9 74.8 77.0
+Falcon-Mamba-7B LoRA 3.7M 6.5 30.5 40.6 14.9 56.4 42.2 31.8 38.4 32.7
+Lily (Δ + $\Delta\text{ + in}$) 3.7M 44.9 66.8 65.0 10.5 57.1 78.7 64.6 68.2 57.0
+Lily (in) 3.3M 60.2 61.0 67.3 12.9 61.5 80.0 67.5 65.8 59.5
 Table 2: Commonsense reasoning results for LLaMA3-8B across eight tasks.
 represents results taken from Liu et al.
-(2024).
+(2024) and (Wang et al., 2024).
 Bold denotes the highest performance scores for each dataset among different PEFT methods.
-Model PEFT BoolQ PIQA SIQA HellaSwag WinoGrande ARC-e ARC-c OBQA Avg.
-ChatGPT - 73.1 85.4 68.5 78.5 66.1 89.8 79.9 74.8 77.0
-LLaMA3-8B LoRA 70.8 85.2 79.9 91.7 84.3 84.2 71.2 79.0 80.8
-PiSSA 67.1 81.1 77.2 83.6 78.9 77.7 63.2 74.6 75.4
-MiLoRA 68.8 86.7 77.2 92.9 85.6 86.8 75.5 81.8 81.9
-Lily 72.9 85.6 77.8 92.7 83.3 89.7 77.6 82.8 82.8
-Implementation We evaluate Lily on commonsense reasoning with LLMs.
-Regarding the implementation, we utilize LLaMA3-8B (AI@Meta, 2024) and Falcon-Mamba-7B (Zuo et al., 2024) as backbones.
-LLaMA3 is a near-SOTA open-source large language model, while Falcon-Mamba is the latest and only open-source large language model based on the Mamba architecture.
-Using these models allows us to validate the effectiveness of Lily for fine-tuning LLMs and whether this effectiveness can be transferred to architectures beyond Transformers (Mamba, in this case).
+Model PEFT Params BoolQ PIQA SIQA HellaSwag WinoGrande ARC-e ARC-c OBQA Avg.
+ChatGPT - - 73.1 85.4 68.5 78.5 66.1 89.8 79.9 74.8 77.0
+LLaMA3-8B LoRA 56M 70.8 85.2 79.9 91.7 84.3 84.2 71.2 79.0 80.8
+PiSSA 83.8M 67.1 81.1 77.2 83.6 78.9 77.7 63.2 74.6 75.4
+MiLoRA 56.6M 68.8 86.7 77.2 92.9 85.6 86.8 75.5 81.8 81.9
+Lily 1.2M 72.9 85.6 77.8 92.7 83.3 89.7 77.6 82.8 82.8
+Implementation: We evaluate Lily on commonsense reasoning with LLMs.
+For the implementation, we utilize LLaMA3-8B (AI@Meta, 2024) and Falcon-Mamba-7B (Zuo et al., 2024) as backbones.
+LLaMA3 is a near-SOTA open-source large language model, while Falcon-Mamba is an open-sourced large language model based on the Mamba architecture.
+Using these models allows us to validate the effectiveness of Lily for fine-tuning LLMs and assess whether this effectiveness can be transferred to architectures beyond Transformers (Mamba, in this case).
 We fine-tune these models on Commonsense170K (Hu et al., 2023) and evaluate the adaptation results on eight multiple-choice problem tasks, including BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-e, ARC-c (Clark et al., 2018), and OBQA (Mihaylov et al., 2018).
@@
 We only compare LoRA for Falcon-Mamba because tailored PEFT methods for Mamba-based LLMs have not yet been proposed, which is beyond the scope of this paper.
-Detailed hyper-parameter settings and datsets information are reported in Appendix B.1.1 and Appendix B.2.1.
-Results We report the accuracy in the Table 2 and Table 1.
-Based on the results, it is evident that Lily performs the best out of the compared PEFT methods.
-Lily surpasses LoRA by a significant margin on Falcon-Mamba, and on LLaMA3, it outperforms LoRA and MiLoRA.
-This indicates Lily’s superior adaptation capability and parameter efficiency dealing with commonsense reasoning tasks.
-Additionally, while the performance on Falcon-Mamba is notably lower than the baseline and LLaMA3, we believe this is due to the model’s limitations rather than Lily’s, as Lily still significantly outperforms LoRA on Falcon-Mamba and demonstrates great performance on LLaMA3.
-This sheds light on the current state of Mamba-based LLMs, showing that they generally have inferior performance compared to Transformer-based LLMs like ChatGPT and LLaMA on many tasks.
+Detailed hyper-parameter settings and dataset information are reported in Appendix B.1.1 and Appendix B.2.1.
+Results We report the accuracy in Tables 2 and 1.
+Based on these results, it can be observed that Lily outperforms the other compared PEFT methods with a smaller parameters budget.
+Specifically, Lily surpasses LoRA by a significant margin on Falcon-Mamba and, on LLaMA3, outperforms both LoRA and MiLoRA.
+This demonstrates Lily’s superior adaptation capability and parameter efficiency in handling commonsense reasoning tasks.
+Additionally, although performance on Falcon-Mamba is notably lower than that of the baseline and LLaMA3, we believe this discrepancy stems from the inherent limitations of the model rather than any deficiency in Lily, as Lily still significantly outperforms LoRA on Falcon-Mamba while demonstrating robust performance on LLaMA3.
+These findings also highlight that current state of Mamba-based LLMs generally exhibits inferior performance compared to Transformer-based LLMs such as ChatGPT and LLaMA on many tasks.
 4.2 Natural Language Understanding
-Table 3: Various fine-tuning methods applied to RoBERTa Base are evaluated on 6 datasets from the GLUE benchmark.
-We present the Matthew’s correlation coefficient (MCC) for CoLA, Pearson correlation coefficient (PCC) for STS-B, and accuracy (Acc.) for the remaining tasks.
-The highest performance for each dataset is highlighted in bold, with all metrics favoring higher values across the 6 datasets.
+Table 3: Various fine-tuning methods applied to RoBERTa Base and RoBERTa Large are evaluated on six datasets from the GLUE benchmark.
+We present the Matthews correlation coefficient (MCC) for CoLA, the Pearson correlation coefficient (PCC) for STS-B, and accuracy (Acc.) for the remaining tasks.
+The highest performance for each dataset is highlighted in bold, with all metrics favoring higher values across the six datasets.
 Model & Method
@@
 $\rm{RoB_{base}}$($\text{Adpt}^{\text{D}}$) 0.9M 94.7 88.4 62.6 93.0 75.9 90.3 84.2
-$\rm{RoB_{base}}$(LoRA) 0.3M 95.1 89.7 63.4 93.3 78.4 91.5 85.2
+$\rm{RoB_{base}}$(LoRA) 0.3M 94.8 89.8 63.3 92.9 78.2 91.5 85.1
 $\rm{RoB_{base}}$(AdaLoRA) 0.3M 94.5 88.7 62.0 93.1 81.0 90.5 85.0
 $\rm{RoB_{base}}$(DyLoRA) 0.3M 94.3 89.5 61.1 92.2 78.7 91.1 84.5
+$\rm{RoB_{base}}$(FLoRA) 0.3M 91.2 85.8 65.4 92.3 65.0 87.6 81.2
 $\rm{RoB_{base}}$(Lily) 0.3M 95.0 90.2 66.0 92.5 81.6 90.8 86.0
+$\rm{RoB_{large}}$(FF) 356M 96.4 90.9 68 94.7 86.6 92.4 88.2
+$\rm{RoB_{large}}$($\text{Adpt}^{\text{H}}$) 0.8M 96.3 87.7 66.3 94.7 72.9 91.5 84.9
+$\rm{RoB_{large}}$(LoRA) 0.8M 96.2 90.2 68.2 94.8 85.2 92.3 87.8
+$\rm{RoB_{large}}$(Lily) 0.5M 95.6 90.9 68.4 94.8 88.4 91.9 88.4
 Implementation We evaluate Lily on natural language understanding (NLU) tasks.
-For implementation, we use RoBERTa Base (Liu et al., 2019) as the backbone and fine-tune it on tasks from GLUE benchmark (General Language Understanding Evaluation (Wang et al., 2018)), consisting of multiple NLU tasks including single-sentence classification tasks, similarity and paraphrase tasks and natural language inference tasks.
-We compare Lily against several competitive PEFT methods, including BitFit (Zaken et al., 2021), Adapter-Tuning (Rücklé et al., 2020) (Houlsby et al., 2019a) (Lin et al., 2020) (Pfeiffer et al., 2020a), LoRA (Hu et al., 2021), DyLoRA (Valipour et al., 2022) and AdaLoRA (Zhang et al., 2023).
+For the implementation, we use RoBERTa Base (Liu et al., 2019) and RoBERTa Large as the backbones and fine-tune them on tasks from the GLUE benchmark (General Language Understanding Evaluation (Wang et al., 2018)), which consists of multiple NLU tasks, including single-sentence classification, similarity and paraphrase, and natural language inference tasks.
+We compare Lily against several competitive PEFT methods, including BitFit (Zaken et al., 2021), Adapter-Tuning (Rücklé et al., 2020; Houlsby et al., 2019b; Lin et al., 2020; Pfeiffer et al., 2020b), LoRA (Hu et al., 2021), DyLoRA (Valipour et al., 2022), FLoRA (Hao et al., 2024), and AdaLoRA (Zhang et al., 2023).
 Additionally, we utilize full fine-tuning (FFT) as the baseline.
-Specific hyper-parameters and datasets information are provide in Appendix B.1.2 and B.2.2.
-Results The results are shown in Table.
-3, from which we can clearly observe that Lily surpass all of the compared PEFT methods by a significant margin, demonstrating its capability of tackling NLU tasks.
-Among the 6 given tasks, Lily surpasses FFT in 4 of them, showcasing its strong approximation ability with compatible parameter-efficiency.
-Figure 2: Results of subject-driven generation. Lily’s results align better with prompts, featuring more accurate color, environment, and shape.
+Specific hyperparameters and dataset information are provided in Appendix B.1.2 and B.2.2.
+Results The results are shown in Table 3.
+From the table, we can clearly observe that Lily surpasses all the compared PEFT methods by a significant margin, demonstrating its ability to tackle NLU tasks.
+Among the six tasks, Lily surpasses FFT on four of them when using RoBERTa Base and RoBERTa Large, showcasing its strong approximation ability and high parameter efficiency.
+Figure 2: Qualitative results of subject-driven generation. Lily’s results align better with prompts, featuring more accurate color, environment, and shape.
 4.3 Subject-driven Image Generation
-Implementation We conduct experiments on fine-tuning text-to-image diffusion models for the subject-driven generation task (Ruiz et al., 2023).
-For backbone, we use SDXL and we fine-tune it using LoRA and Lily.
-We first fine-tune the model with images associated with text prompts (e.g., A photo of a [v] duck toy), in which a unique identifier is provided.
-After that, text prompts containing the identifier could be used to generate customized images.
+Implementation We conduct experiments on fine-tuning text-to-image diffusion models for the subject-driven generation task.
+As the backbone, we use SDXL and fine-tune it using both LoRA and Lily.
+First, we fine-tune the model on images paired with text prompts (e.g., “A photo of a [v] duck toy”), each of which includes a unique identifier.
+Afterward, text prompts containing the identifier are used to generate customized images.
 Results The results are presented in Fig.
@@
 (2024b) and Wu et al.
-(2024), from which we can observe that images generated by Lily generally align better with the text prompts.
-For instance, when asked to generate a duck toy floating on top of water, Lily’s image accurately depicts the designated environment, whereas LoRA’s does not.
-Additionally, when asked to generate a wolf plushie in snow, Lily precisely depicts the snow around the wolf, while LoRA fails to do so.
-These observations demonstrate Lily’s excellent ability in the domain of text-to-image generation with more expressive adaptation.
-More generated results are in Appendix I.
+(2024).
+From these results, we observe that the images generated by Lily generally align better with the text prompts.
+For instance, when asked to generate an image of a duck toy floating on water, Lily’s output accurately depicts the designated environment, whereas LoRA’s does not.
+Additionally, when asked to generate an image of a wolf plushie in the snow, Lily precisely captures the snow around the wolf, while LoRA fails to do so.
+These observations demonstrate Lily’s excellent performance in text-to-image generation with more expressive adaptation.
+Additional generated results are provided in Appendix I.
 4.4 Visual Adaptation Benchmark
@@
 * denotes linear probing results from Tu et al.
-(2023b).
+(2023a).
 For fair comparison, we also use ViT-B pre-trained on ImageNet-1K.
@@
 Lily-L 0.196 72.3 57.8 89.4 66.2 87.8 90.5 88.1 40.5 84.1 94.3 81.3 75.1 81.6 57.8 46.5 81.0 82.9 55.2 32.1 34.8
-Implementation We assess Lily on the Visual Task Adaptation Benchmark (VTAB-1K Zhai et al.
-(2019)), a suite of 19 visual tasks spanning diverse domains and semantics, to test its general visual adaptation capability.
-Tasks are categorized into Natural, Specialized, and Structured, all formulated as classification problems for consistent model evaluation.
-We conduct two sets of experiments: one focusing on the adaptation effectiveness on Vision Transformer (ViT (Dosovitskiy et al., 2020)) and the other on Vision Mamba (Vim (Zhu et al., 2024)), demonstrating Lily’s architecture-agnostic capabilities.
-For ViT, we use ViT-B pre-trained on ImageNet-21K (Deng et al., 2009), and for Vim, Vim-s pre-trained on ImageNet-1K.
-To fairly compare ViT and Vim architectures, we implement LoRA (Hu et al., 2021) and AdaptFormer (Chen et al., 2022) on ViT-B pre-trained on ImageNet-1K.
-In ViT experiments, we compare Lily with LoRA, AdaptFormer, FourierFT (Gao et al., 2024b), and MoRA (Jiang et al., 2024); in Vim experiments, we focus on contrasting architectures difference, therefore only using LoRA as the baseline.
-All experiments include FFT and linear probing as baselines.
-For Vim, we implement two versions: Lily-S (Small) and Lily-L (Large) of Lily, with different hyperparameter settings to either reduce the parameter count (Lily-S) or maximize performance (Lily-L).
-For Lily on ViT, the reported results are obtained from adapting both the self attention and the MLP module in Transformer.
-For the performance w.r.t the fine-tuned module, we conduct additional experiments in Appendix D.
-Detailed experimental settings and datasets information are provide in Appendix B.1.3 and B.2.3.
-Results Results are shown in Table 4 and Table 5.
-For ViT, Lily significantly outperforms all compared PEFT methods with improved parameter efficiency.
-For Vim, results on ViT generally surpass those on Vim.
-For instance, LoRA on ViT performs better than LoRA on Vim.
-We argue that this is due to differences in architecture designs and general model sizes.
-However, Lily’s strong adaptation performance allows it to match or exceed PEFT methods on ViT and significantly outperform LoRA on Vim (Lily-S and Lily-L surpass LoRA by a significant margin).
+[…]
+Results The results are shown in Tables 4 and 5.
+For ViT, Lily significantly outperforms all compared PEFT methods while also offering improved parameter efficiency.
+In contrast, the performance on the Vim backbone is generally lower than that on ViT; for instance, LoRA on ViT performs better than LoRA on Vim.
+We argue that this difference is due to variations in architecture design and overall model size.
+However, Lily’s strong adaptation performance allows it to match or exceed the performance of other PEFT methods on ViT and to significantly outperform LoRA on Vim (with both Lily-S and Lily-L surpassing LoRA by a significant margin).
 This demonstrates Lily’s architecture-agnostic capability, highlighting its potential across various model architectures.
-In general, Lily has achieved great visual adaptation capability with an advantage of being architecture-agnostic and enjoying excellent parameter-efficiency.
-Figure 3: Attention maps of Lily and LoRA.
-The input images for the example here are taken from Caltech101 datasets from VTAB-1K benchmark.
-It can be observed that features from a certain layer have more similarity to those in other layers in Lily than in LoRA.
-Figure 4: Feature difference measured in absolute distance for each element.
-We compare Lily and LoRA in terms of the difference between features from different layers.
-In this example image taken from Caltech101, we visualize the feature difference between layers 6 and 1, as well as between layers 6 and 9.
+Overall, Lily achieves excellent visual adaptation performance while maintaining architecture-agnosticity and high parameter efficiency.
 4.5 Understanding Lily
-From a Feature Merging Perspective.
-Apart from being a better gradient projector than LoRA, Lily also enables comprehensive information access across layers.
-Lily enables access to information or features from all other layers when adapting a target module at a specific layer thanks to the inter-connectivity of the adapters.
-We aim to understand how Lily achieves this comprehensive information access from the perspective of visual tasks as shown in Fig.
+4.5.1 Does It Have High-Rank Weight Updates?
+The interconnected and asymmetric structure of Lily enables a flexible allocation of the parameter budget, thereby allowing weight updates with higher ranks across all layers.
+To validate this claim, we provide an empirical analysis, as shown in Fig.
 3.
-We can observe that, in Lily, the distinctness of the attention maps between layers is not as pronounced as in LoRA.
-This validates Lily’s ability to enable all-level information access, since adaptation at each layer takes into account features from other layers.
-Additionally, we specifically visualize the actual feature differences between different layers in Fig.
-4.
-We observe that Lily has more points with low feature differences (blue color) than LoRA, indicating that the distinctness of features between layers in Lily is generally lower than in LoRA.
-This further demonstrates Lily’s ability to enable comprehensive information access.
-Although we enable all-level information access, what prevents the features from becoming completely identical is the selectivity introduced by Lily, which we specify in the following section.
-Figure 5: Impact of attention granularity (i.e., the choice of how many LPs and HPs) on the performance. We choose 12 out of 19 tasks from VTAB-1K for a comprehensive understanding.
-Figure 6: Visualization of accumulated assigned weight for HP experts by a router across various layers. Example here uses layer of index 2, 13 and 22 to represent shallow, middle and deep layers.
-What’s the Influence of Attention Granularity?
-The number of experts in the model-wide HP module can be freely set, and the number of LPs can also be flexibly set by sharing across the same level of layers introduced in Appendix A.1.
+Specifically, we run four tasks from the NLU experiment and measure the rank of the weight updates for the query transformation matrix $W_{q}$ in the first three layers.
+We use a small number of matrices $A$ and $B$ (2 or 3) with a rank of 32 to match the parameter count of LoRA, which uses adapters with a rank set to 8.
+Specific hyperparameter settings can be found in Appendix B.1.2.
+From the results, we observe that the rank of the weight updates from Lily is generally notably larger than that of LoRA when using a similar number of parameters.
+Meanwhile, the weight updates from Lily still exhibit a higher rank compared to those of LoRA even when using only $16.7\%$ of LoRA’s parameters.
+This empirical analysis essentially validates our claim that Lily achieves high-rank updates with the same parameter budget.
+We attribute this to the model-wide sharing and the cross-layer asymmetric design, which facilitate a flexible allocation of the parameter budget.
+Figure 3: Actual rank of the weight updates.
+The weight updates are of shape $768\times 768$.
+We run 20 epochs for COLA, MRPC, and STS-B, and 3 epochs for SST-2.
+It can be easily observed that the weight updates from Lily have notably higher rank than those from LoRA.
+Note that the reported rank is computed from accumulated weight updates over multiple epochs.
+Figure 4: Visualization of accumulated assigned weight for $B$ experts by a router across various layers.
+Example here uses layer of index 2, 13 and 22 to represent shallow, middle and deep layers.
+The reported values are based on the accumulated router outputs over multiple epochs.
+Figure 5: Impact of attention granularity (i.e., the choice of how many $A$s and $B$s) on the performance. We choose 12 out of 19 tasks from VTAB-1K for a comprehensive understanding.
+Figure 6: Hardware efficiency of Lily compared to LoRA.
+We run 10 epochs for COLA.
+We report the training time and memory consumption.
+It can be observed that Lily generally performs on par with LoRA in terms of hardware efficiency.
+4.5.2 What’s the Influence of Adapter Granularity?
+The number of experts in the model-wide $B$ module can be freely set, and the number of $A$s can also be flexibly determined by sharing across the same level of layers, as introduced in Appendix A.1.
 Therefore, we analyze the impact of these choices on performance.
-We denote the number of LP experts and HP experts as ne_1 and ne_2, respectively.
-For simplicity, we make them identical in the experiments, denoted as ne.
-We refer to the number of layers each expert attends to as attention granularity.
-As the value of ne increases, the attention granularity becomes finer.
+We denote the number of $A$ experts and $B$ experts as ne_1 and ne_2, respectively.
+For simplicity, we set them equal in the experiments and denote this common value as ne.
+We refer to the number of layers each expert attends to as adapter granularity.
+As the value of ne increases, the adapter granularity becomes finer.
 As shown in Fig.
 5, the results from the VTAB-1K benchmark indicate different patterns.
-For instance, on the DTD dataset, the best performance is achieved when ne is 4, while on sNORB-Azim, performance increases with the increase in ne.
-Increasing ne leads to more parameters and finer attention granularity.
-However, finer attention granularity does not necessarily lead to better overall performance.
-For example, on Resisc45, DTD, Cifar100, sNORB-Ele, dsPr-LoC, Flowers102, and EuroSAT, the negative impact of increasingly finer attention granularity eventually outweighs the benefits of increased parameters, leading to a decrease in overall performance.
-In other tasks, different patterns may occur because the positive effect of attention granularity on performance is consistently strong, or its negative effect is not enough to offset the benefits of increased parameters, resulting in a generally increasing performance with ne.
-This phenomenon provides an important insight: for most tasks, simply increasing parameters may not lead to better performance.
-Instead, only when attention granularity and the number of parameters reach a good tradeoff can we achieve the best performance.
-Does Lily Exhibit Selectivity?
-Lily uses routers to assign varying weights to different HP experts, thereby achieving selective information combination.
+For instance, on the DTD dataset, the best performance is achieved when ne is 4, while on sNORB-Azim, performance increases as ne increases.
+Increasing ne leads to more parameters and finer adapter granularity; however, finer adapter granularity does not necessarily translate to better overall performance.
+For example, on Resisc45, DTD, Cifar100, sNORB-Ele, dsPr-LoC, Flowers102, and EuroSAT, the negative impact of increasingly finer adapter granularity eventually outweighs the benefits of the additional parameters, leading to a decrease in overall performance.
+In other tasks, different patterns may occur because the positive effect of adapter granularity on performance is consistently strong, or its negative effect is insufficient to offset the benefits of increased parameters, resulting in generally improved performance with higher ne.
+This phenomenon provides an important insight: for most tasks, simply increasing the number of parameters may not lead to better performance.
+Instead, only when adapter granularity and the number of parameters reach an optimal trade-off can we achieve the best performance.
+4.5.3 Does It Exhibit Selectivity?
+Lily uses routers to assign varying weights to different $B$ experts, thereby achieving selective information combination.
 We illustrate this selectivity in Fig.
-6.
-We use a setup with three HP experts and select three layer levels (1, 13, 22) to calculate the total weight assigned to each expert.
-The results reveal a clear selectivity: for different layers, the router assigns significantly different weights to different HP experts.
-For instance, on Cifar100, the middle layer is predominantly dominated by HP 2, whereas the deep layer is primarily dominated by HP 1 and HP 2.
-In contrast, on Retinopathy, both the middle and deep layers are dominated by HP 3.
+4.
+We use a setup with three $B$ experts and select three layer levels (1, 13, 22) to calculate the total weight assigned to each expert.
+The results reveal clear selectivity: for different layers, the router assigns significantly different weights to the $B$ experts.
+For instance, on Cifar100, the middle layer is predominantly dominated by $B$ 2, whereas the deep layer is primarily dominated by $B$ 1 and $B$ 2.
+In contrast, on Retinopathy, both the middle and deep layers are dominated by $B$ 3.
 This selectivity ensures that, even when different layers share information, the inherent differences between layers are still taken into account, making the adaptation more flexible and comprehensive.
+4.5.4 What’s the Hardware Efficiency?
+The dynamics of Lily obviously introduce complexity to the design of LoRA.
+In this section, we analyze how this affects the hardware efficiency of Lily compared to LoRA.
+We use the COLA task from the NLU experiments with RoBERTa-Base and run for 10 epochs.
+Additionally, we also report the runtime and GPU memory consumption in the Falcon-Mamba experiment.
+The results are shown in Fig.
+6, from which we can observe that the hardware efficiency of Lily is comparable to LoRA.
+Specifically, Lily slightly underperforms LoRA in the NLU experiment but performs on par with LoRA in the LLM experiment.
+In general, the introduced complexity of Lily does not prevent it from being an more effective PEFT method that is also hardware-friendly.
 5 Conclusion
-In this paper, we propose low-rank interconnected adaptation (Lily), a novel framework for efficient fine-tuning via inter-connectivity of adapters.
+In this paper, we propose Low-Rank Interconnected Adaptation (Lily), a novel framework for efficient fine-tuning via the interconnectivity of adapters.
 Lily enables each layer to access information from others during adaptation through a hierarchical structure.
-Additionally, it successfully overcome the low-rank update limitation of LoRA, enabling high-rank update and therefore better adaptation capability.
-Our approach consistently improves performance across various modalities, model sizes, and architectures, surpassing existing methods with enhanced efficiency.
+Additionally, it successfully overcomes the low-rank update limitation of LoRA, enabling high-rank updates and, therefore, better adaptation capability under the same parameter budget.
+Our approach consistently improves performance across various modalities, model sizes, and architectures, surpassing existing methods while maintaining enhanced efficiency.
 In summary, Lily’s versatility and efficiency make it a promising approach for a wide range of applications.
+6 Limitations
+Although Lily has been experimentally evaluated in a wide range of scenarios, we have not explored all possible applications where PEFT could be used.
+These potential areas are left as directions for future work.
+7 Ethics Statement
+This work is an improvement upon LoRA. However, it could potentially be used for fine-tuning diffusion models or large language models (LLMs) for generating malicious content.
 Appendix
@@
 Within the overall framework of Lily, we delve into specific implementation details and model design insights.
-First, we have established the relationship between LPs and HP: LP is confined to specific levels of layers, capturing features that enable the router to selectively assign weights to the HP experts.
-In contrast, HP is a model-wide module comprising multiple experts, each of which contains information from a particular level of layers.
-We highlight several key aspects which are not heavily discussed in the methodology section:
-A.1.1 Number of LPs
-Since LP is limited to specific layers, the simplest approach would be to place an LP at each layer of the module to be adapted (e.g., the query transformation in MHSA).
-However, this setup may not be necessary: the importance of each layer varies, and many layers have significantly lower importance than others [Zhang et al., 2023].
-To achieve greater parameter efficiency, we can set up fewer LPs, with each LP focusing on a level of layers rather than a single layer.
-For example, an LP can focus on shallow layers (e.g., layers 0, 1, 2, etc.) or deep layers.
-To enable a single LP to handle multiple layers, we can share an LP across multiple layers.
-By doing so, we eliminate the redundancy of having an LP at each layer, reduce the number of parameters, and increase efficiency.
-This is exactly the strategy adopted by most of the experiments.
-A.1.2 Number of HP Experts
-Regarding HP, the number of experts can be arbitrarily set, enabling more flexible configurations.
-In our experiments, for the sake of simplicity, we set the number of HP experts equal to the number of LPs, thereby equating the granularity of LP and HP.
+First, we establish the relationship between $A$ and $B$: $A$ is confined to specific levels of layers, capturing features that enable the router to selectively assign weights to the $B$ experts.
+In contrast, $B$ is a model-wide module comprising multiple experts, each of which contains information from a particular level of layers.
+We highlight several key aspects that are not heavily discussed in the methodology section:
+A.1.1 Number of $A$s
+Since $A$ is limited to specific layers, the simplest approach would be to place an $A$ at each layer of the module to be adapted (e.g., the query transformation in MHSA).
+However, this setup may not be necessary, as the importance of each layer varies, and many layers have significantly lower importance than others (Zhang et al., 2023).
+To achieve greater parameter efficiency, we can use fewer $A$s, with each $A$ focusing on a level of layers rather than a single layer.
+For example, an $A$ can focus on shallow layers (e.g., layers 0, 1, 2, etc.) or deep layers.
+To enable a single $A$ to handle multiple layers, we can share an $A$ across multiple layers.
+By doing so, we eliminate the redundancy of placing an $A$ at each layer, reduce the number of parameters, and improve efficiency.
+This is exactly the strategy adopted in most of the experiments.
+A.1.2 Number of $B$ Experts
+Regarding $B$, the number of experts can be set arbitrarily, allowing for more flexible configurations.
+In our experiments, for the sake of simplicity, we set the number of $B$ experts equal to the number of $A$s, thereby equating the granularity of $A$ and $B$.
 A.1.3 Routers Setup
-There are also different settings that can be employed for the router.
-First, we can bind the router to HP, resulting in only one router per model.
-However, since the number of parameters in the router is relatively small, having only one router per model may not lead to significant selectivity.
-Therefore, we can also bind the router to LP, configuring a separate router for each LP.
-Most of our experiments use the latter setup, but in the vision experiments on Vim, we use the single-router and no-lp-sharing setup to verify its effectiveness.
-From the results, we can see that this setup also performs well.
-As future work, we can verify the effectiveness of using the latter setup on Vim, which may potentially lead to superior performance.
+There are multiple possible configurations for the router.
+First, we can bind the router to $B$, resulting in only one router per model.
+However, since the number of parameters in the router is relatively small, having only one router per model may not provide significant selectivity.
+Therefore, we can also bind the router to $A$, configuring a separate router for each $A$.
+Most of our experiments use the latter setup.
+However, in the vision experiments on Vim, we adopt the single-router and no-lp-sharing setup to evaluate its effectiveness.
+The results indicate that this setup also performs well.
+As future work, we can further verify the effectiveness of the latter setup on Vim, which may potentially lead to superior performance.
 A.1.4 Hyperparameters
 We detail the hyperparameters used in Lily.
-Specifically, we use lily_r to represent the hidden-dimension of the projectors: LPs and HPs.
+Specifically, we use Lily_r to represent the hidden dimension of the projectors: $A$s and $B$s.
 It serves the same function as $r$ in LoRA.
-We use lily_s to represent the scaling factor used by Lily.
-It is mostly searched within the range of {0.01, 0.1, 1.0, 10.0, 100.0}.
-We use ne_1 to represent the number of LPs used in the model.
-Since the LPs can be shared as discussed in the previous section, ne_1 does not need to equal the number of layers in the model.
-We use ne_2 to represent the number of HP experts in the model-wide HP module.
-In our experiments, we set ne_1 = ne_2 to improve parameter-efficiency and simplicity.
+We use Lily_s to represent the scaling factor used by Lily, which is primarily searched within the range {0.01, 0.1, 1.0, 10.0, 100.0}.
+We use ne_1 to denote the number of $A$s used in the model.
+Since $A$s can be shared, as discussed in the previous section, ne_1 does not necessarily equal the number of layers in the model.
+Similarly, we use ne_2 to represent the number of $B$ experts in the model-wide $B$ module.
+In our experiments, we set ne_1 = ne_2 to enhance parameter efficiency and maintain simplicity.
 A.1.5 Design Intuition
-Lily employs a hierarchical structure to enable updates with higher-ranks than LoRA.
-However, simply equally connecting all the HPs to the LPs can not achieve the best performance.
-From the perspective of feature and information utilization across layers, simply aggregating all HPs for an LP ignores the distinctness of the features from current layers.
-Meanwhile, it reduced the variability of the combinations of gradient projection matrices ($S_{i}$ and $C_{i,j}$ are constants now), making the rank of the weight update higher than that of LoRA (since multiple distinct random matrices are used), but still not high enough for the best performance because of the lack of variability during combination.
-Therefore, we introduce selectivity into the inter-connectivity as well as discussed below, making the combination of HPs data-dependent so that each $S_{i}$ is unique across the time-steps, enabling updates with even higher ranks.
-The approach has similarity to Hao et al.
-[2024], where random matrix is constantly resampled to ensure updates with higher ranks.
-We further conduct an analysis in Appendix G.
+Lily employs a hierarchical structure to enable updates with higher ranks than LoRA.
+However, simply connecting all $B$s equally to $A$s does not yield the best performance.
+From the perspective of feature and information utilization across layers, merely aggregating all $B$s for an $A$ ignores the distinctiveness of features from the current layers.
+Meanwhile, this approach reduces the variability in the combinations of gradient projection matrices (since $S_{i}$ and $C_{i,j}$ become constants), making the rank of the weight update higher than that of LoRA (as multiple distinct random matrices are used), but still not high enough for optimal performance due to the lack of variability in the combination process.
+To address this, we introduce selectivity into the interconnectivity, as discussed below, making the combination of $B$s data-dependent.
+This ensures that each $S_{i}$ is unique across time steps, enabling updates with even higher ranks.
+This approach is similar to that of Hao et al.
+(2024), where a random matrix is constantly resampled to maintain high-rank updates.
+We further analyze this in Appendix G.
 A.2 Efficient Implementation for Weighted Combination
-One intuitive implementation of the weighted combination in Lily is Eq.
-17, from which we could observe that it perform $N_{e}$ times of matrix multiplication, $N_{e}$ times of scalar multiplication and $N_{e}$ times of matrix addition.
-Therefore, despite its intuitive nature, the computational burden of this approach is quite formidable.
+A straightforward implementation of the weighted combination in Lily is to pass the inputs through all the experts and then sum the results.
+This approach requires $N_{e}$ matrix multiplications, $N_{e}$ scalar multiplications, and $N_{e}$ matrix additions.
+Despite its intuitive nature, the computational burden of this method is quite substantial.
 However, Eq.
-18 which is adopted in Lily only utilize $N_{e}$ times of scalar multiplication, $N_{e}$ times of matrix addition and 1 time of matrix multiplication.
-This saves roughly $N_{e}$ times of matrix multiplication, which can be significant as the size of the model and number of adaptation targets increases.
-For a $x^{\prime}$ size of $\mathbb{R}^{N\times d}$ and a $P_{H}\in\mathbb{R}^{d\times C}$, the floating-point operation (FLOPs) of these two implementation are:
-FLOPs $=\sum_{1}^{N_{e}}(2NdC)+\sum_{1}^{N_{e}}(dC)+\sum_{1}^{N_{e}}(NC)$ (21)
-$=N_{e}\times(2NdC+dC+NC)$ (Intuitive)
+3, which is adopted in Lily, requires only $N_{e}$ scalar multiplications, $N_{e}$ matrix additions, and a single matrix multiplication.
+This optimization eliminates approximately $N_{e}$ matrix multiplications, which can significantly reduce computational costs as the model size and the number of adaptation targets increase.
+For an input $x^{\prime}$ of size $\mathbb{R}^{N\times d}$ and a projection matrix $P_{H}\in\mathbb{R}^{d\times C}$, the floating-point operations (FLOPs) of these two implementations are:
+FLOPs $=\sum_{i=1}^{N_{e}}(2NdC)+\sum_{i=1}^{N_{e}}(dC)+\sum_{i=1}^{N_{e}}(NC)$ (5)
+$=N_{e}\times(2NdC+dC+NC),$
 FLOPs $=2\sum_{i=1}^{N_{e}}(dC)+2NdC$
-$=2dC\times(N+N_{e})$ (Lily)
-from which we can easily observe that the approach adopted by Lily requires less computation and therefore provides more speed and efficiency during the fine-tuning process.
-Under the setting of $N=1024,d=16,C=768,N_{e}=4$, the FLOPs of the intuitive approach would be $0.104$ GFLOPs while in Lily it is merely $0.025$ GFLOPs, which could potentially lead to a 4X increase in speed.
+$=2dC\times(N+N_{e}),$
+From this, we can easily observe that the approach adopted by Lily requires fewer computations, thereby improving both speed and efficiency during the fine-tuning process.
+Under the setting of $N=1024,d=16,C=768,N_{e}=4$, the FLOPs for the intuitive approach amount to $0.104$ GFLOPs, whereas for Lily, it is merely $0.025$ GFLOPs, potentially leading to a 4×4$\times$4 × speedup.
 A.3 Actual Implementation of Lily
-Figure 7: Implementation of Lily in VTAB-1K benchmark.
+Figure 7: Implementation of Lily in the VTAB-1K benchmark.
 We present the actual implementation of Lily in Fig.
 7.
-For the example here, we choose the implementation from visual adaptation tasks (i.e., VTAB-1K benchmark).
-For LLM, the implementation is a bit more complicated because of modifications to the huggingface PEFT library [Mangrulkar et al., 2022], but the fundamental adaptation process is the same.
-Specifically, given an input, we first use the corresponding LP of the current layer to project it to a low-dimensional representation.
-After that, we use the low-dimensional representation to selectively assign weights for the HP experts.
-Once we obtain all the weights for the experts, we set out to combine these HP experts accordingly, as discussed in Appendix A.2.
-After the weighted combination, we use the obtained combined HP to project the low-dimensional representation to high-dimension, therefore acquiring the extra knowledge gained through adaptation.
+In this example, we showcase its implementation for visual adaptation tasks, specifically in the VTAB-1K benchmark.
+For large language models (LLMs), the implementation is slightly more complex due to modifications to the Hugging Face PEFT library (Mangrulkar et al., 2022), but the fundamental adaptation process remains the same.
+Specifically, given an input, we first use the corresponding $A$ of the current layer to project it into a low-dimensional representation.
+This low-dimensional representation is then used to selectively assign weights to the $B$ experts.
+Once the weights for all experts are obtained, we proceed to combine these $B$ experts accordingly, as discussed in Appendix A.2.
+After the weighted combination, the combined $B$ is used to project the low-dimensional representation back into a high-dimensional space, thereby incorporating the additional knowledge gained through adaptation.
 Appendix B Experimental Settings
-B.1 Hyper-parameters
-A detailed description of the hyper-parameters used in Lily is in Appendix A.1.
+B.1 Hyperparameters
+A detailed description of the hyperparameters used in Lily is provided in Appendix A.1.
 B.1.1 Commonsense Reasoning
-The hyper-parameters used in commonsense reasoning experiments for MiLoRA, PiSSA are provided in Table 7, 6.
-Settings for Lily and LoRA using Falcon-Mamba as the backbone are provided in Table.
-9 and 8.
-It can be noticed that Lily achieves the best performance by merely adapting the multi-head self attention module (MHSA) in LLaMA3-8B, while other compared methods adapt all the modules including MLP.
-Meanwhile, Lily employs the least amount of parameters, showcasing its excellent adaptation at low parameter-budget scenarios.
+The hyperparameters used in commonsense reasoning experiments for MiLoRA and PiSSA are provided in Tables 7 and 6.
+The settings for Lily and LoRA using Falcon-Mamba as the backbone are presented in Tables 9 and 8.
+[…]
 Table 6: Hyperparameter configuration from the MiLoRA paper.
@@
 Epochs 3
-Placement query, key, value, MLP up, MLP down
+[…]
 Table 7: Hyperparameter configuration from the PiSSA paper.
@@
 Epochs 1
-Placement query, key, value, output, gate, MLP up, MLP down
+[…]
 Table 8: Hyperparameter configuration for LoRA using Falcon-Mamba as backbone.
@@
 B.1.2 Natural Language Understanding
-Specific hyper-parameter settings of Lily on GLUE benchmark are provided in Table 10.
-We fix the learning rate of both the backbone and the head as 5E-3 and tune the scaling factor lily_s $\in\{0.01,0.1,1.0\}$ instead.
-For the rank $r$ we fix it to 32 and the seed to 0.
+The specific hyperparameter settings for Lily on the GLUE benchmark are provided in Table 10.
+We fix the learning rate of both the backbone and the head at $5\times 10^{-3}$ and instead tune the scaling factor Lily_s, where Lily_s $\in\{0.01,0.1,1.0\}$.
+The rank $r$ is fixed at 32, and the random seed is set to 0.
+The baseline results are taken from FourierFT (Gao et al., 2024b).
 Table 11: Hyperparameter configuration for Lily on VTAB-1K benchmark.
@@
 LR Decay cosine cosine
-B.1.3 Visual Adaptation Benchmark
-We provide the hyper-parameter for Lily on VTAB-1K benchmark in Table 11.
-Specifically, we fix the learning rate at 1E-3 with a weight decay of 1E-4.
-For ViT, we tune the scaling factor lily_s $\in\{0.01,0.1,1.0,10.0\}$ to maximize the performance, following Jie et al.
-[2023], Jie and Deng [2023].
+B.1.3 Visual Task Adaptation Benchmark
+We provide the hyperparameters for Lily on the VTAB-1K benchmark in Table 11.
+Specifically, we fix the learning rate at $1\times 10^{-3}$ with a weight decay of $1\times 10^{-4}$.
+For ViT, we tune the scaling factor Lily_s $\in\{0.01,0.1,1.0,10.0\}$ to maximize performance, following Jie et al.
+(2023) and Jie and Deng (2023).
 For Vim, we fix Lily_s to 1.0.
-Additionally, we search for the hyper-parameters ne_1 and ne_2 within the range {2, 3, 4}, as these numbers can divide the number of layers in the ViT model (12 in ViT-B).
-For vim, we use the implementation discussed in section A.1, which does not share LPs across layers.
-Therefore, ne_1 in this setting is fixed to number of layers in Vim (22 in this case), while we search ne_2 in {3,6} and {5,6,17} separately for Lily-S and Lily-L.
-Note that ne is only set for input projection in Vim.
-For delta transformation, we only use a single HP expert to reduce the parameter cost.
-In the experiments of ViT, the rank $r$ is fixed at 16.
-Meanwhile in Vim’s setting, we tune the ranks $r$ for the delta transformation module and the input projection module separately.
-We use $4,4$ and $4,8$ separately for Lily-S and Lily-L.
+Additionally, we search for the hyperparameters ne_1 and ne_2 within the range {2, 3, 4}, as these numbers divide the number of layers in the ViT model (12 in ViT-B).
+For Vim, we use the implementation discussed in Section A.1, which does not share $A$s across layers.
+Therefore, ne_1 in this setting is fixed to the number of layers in Vim (22 in this case), while we search for ne_2 in {3, 6} and {5, 6, 17} separately for Lily-S and Lily-L.
+Note that ne is only set for the input projection in Vim.
+For the delta transformation, we use only a single $B$ expert to reduce the parameter cost.
+In the ViT experiments, the rank $r$ is fixed at 16.
+Meanwhile, in Vim’s setting, we tune the ranks $r$ for the delta transformation module and the input projection module separately.
+We use $(4,4)$ and $(4,8)$ separately for Lily-S and Lily-L.
 B.2 Datasets
@@
 SmallNORB/elevation 12,150 18
-Appendix C Does Sharing LP Results in Inferior Performance?
-As mentioned earlier, we adopted a strategy of sharing the LP across most of our experiments, ensuring that the number of LP and HP experts is consistent.
-This approach offers two benefits: simplicity and enhanced parameter efficiency.
-By sharing the LP, we eliminate the need to set a separate LP for each layer, thereby reducing the overall parameter count.
-Our decision to share the LP is based on the observation of overall redundancy among layers.
-Specifically, different layers have varying levels of importance [Zhang et al., 2023], and some less important layers do not require a dedicated LP.
-By not setting a separate LP for these layers, we avoid introducing extra parameter overhead while having a negligible impact on performance.
-To test that whether sharing LP results in inferior performance, we conduct experiments with no LP sharing on VTAB-1K.
-The results are shown in Table 15, from which we can observe that the best overall performance ($77.3\%$) is the same as that in the LP-sharing setting.
-This indicates that even if we employ one LP for each layer, the performance gain is negligible and many of the parameters are actually redundant.
-However, not sharing LPs results in extra parameter overhead and damages the parameter-efficiency of Lily.
-Therefore, LP-sharing is a great strategy to eliminate redundancy among LPs and boost the parameter-efficiency of Lily.
+Appendix C Does Sharing $A$s Result in Inferior Performance?
+As mentioned earlier, we adopted a strategy of sharing the $A$ across most of our experiments, ensuring that the number of $A$s and $B$ experts is consistent.
+This approach offers two key benefits: simplicity and enhanced parameter efficiency.
+By sharing the $A$, we eliminate the need to set a separate $A$ for each layer, thereby reducing the overall parameter count.
+Our decision to share the $A$ is based on the observation of overall redundancy among layers.
+Specifically, different layers have varying levels of importance (Zhang et al., 2023), and some less important layers do not require a dedicated $A$.
+By not setting a separate $A$ for these layers, we avoid introducing unnecessary parameter overhead while maintaining negligible impact on performance.
+To test whether sharing $A$ results in inferior performance, we conducted experiments without $A$ sharing on the VTAB-1K benchmark.
+The results, shown in Table 15, indicate that the best overall performance ($77.3\%$) is the same as in the $A$-sharing setting.
+This suggests that even when we employ one $A$ for each layer, the performance gain is negligible, and many of the parameters are, in fact, redundant.
+However, not sharing $A$s leads to additional parameter overhead, which reduces the parameter efficiency of Lily.
+Therefore, $A$-sharing is an effective strategy to eliminate redundancy among $A$s and enhance the parameter efficiency of Lily.
 Appendix D Where to Apply Lily in Transformers?
-Table 15: Performance on VTAB-1K benchmark when applying Lily to various modules in Transformer. The implementation here does not share LP for simplicity (i.e., each layer has one LP).
+Table 15: Performance on VTAB-1K benchmark when applying Lily to various modules in Transformer. The implementation here does not share $A$ for simplicity (i.e., each layer has one $A$).
 Natural Specialized Structured
@@
 kvmlp 77.3 74.0 92.3 72.6 99.3 91.5 89.2 56.7 88.2 95.4 85.3 76.0 84.6 64.9 53.4 81.7 87.5 52.9 36.9 45.2
-PEFT methods have been predominantly explored on the Transformer architecture, which consists of multi-head self-attention (MHSA) and multi-layer perceptron (MLP) as its core modules.
-In this section, we analyze the impact of fine-tuned modules on performance using Lily.
-Specifically, we compare Lily’s performance on the VTAB-1K benchmark under four settings:
+Appendix E Performance Analysis on VTAB-1K Benchmark with Lily on Transformer Modules
+[…]
   • Applying Lily solely to the query and value transformation module in MHSA (denoted as "qv").
-  • Applying Lily solely to the MLP module (denoted as "mlp").
-  • Applying Lily to both the query and value transformation module in MHSA and the MLP module (denoted as "qvmlp").
-  • Applying Lily to both the key and value transformation module in MHSA and the MLP module (denoted as "kvmlp").
+[…]
+[…]
+[…]
 To ensure a fair comparison, we tune the hyperparameters to maintain a similar parameter count across all settings.
-Additionally, to further investigate whether sharing the low-rank projection (LP) affects performance, we do not share the LP in this experiment.
-The results are presented in Table 15.
-We observe that the "kvmlp" setting achieves the best performance, with an average accuracy of $77.3\%$.
-In contrast, adapting only the MHSA module ("qv") yields the worst performance.
-Furthermore, we note that adapting both the MHSA and MLP modules (qvmlp and kvmlp) generally leads to superior results compared to adapting only one specific module (qv and mlp).
-This suggests that both MLP and MHSA play crucial roles in the overall model performance, and adapting both is essential for effective adaptation.
+Additionally, to further investigate whether sharing the low-rank projection ($A$) affects performance, we do not share $A$ in this experiment.
+[…]
 Notably, even when applying Lily solely to the MHSA module, which results in the worst performance among the four settings ($76.9\%$), it still outperforms LoRA by a significant margin ($0.5\%$).
-This underscores the efficiency of Lily, as it uses fewer parameters than LoRA even without LP sharing.
-Appendix E Where to Apply Lily in Mamba?
-Nearly all previous PEFT method studies have been centered around Transformers, while Mamba is a relatively new architecture, so there has been little research on PEFT methods on Mamba.
+This underscores the efficiency of Lily, as it uses fewer parameters than LoRA, even without $A$ sharing.
+Appendix F Where to Apply Lily in Mamba?
+Nearly all previous PEFT method studies have focused on Transformers, while Mamba is a relatively new architecture, and therefore, there has been little research on PEFT methods for Mamba.
 In this section, we briefly analyze the pros and cons of adapting Mamba’s modules.
-In brief, a Mamba block consists of regular linear projection layers and a core component SSM module [Gu and Dao, 2023] [Zhu et al., 2024].
-Specifically in SSM module, Mamba utilize parameters ($\Delta$, $A$, $B$, $C$) to transform an input sequence $x(t)$ to an output sequence $y(t)$ using a hidden state $h(t)$.
+A Mamba block consists of regular linear projection layers and a core component, the SSM module (Gu and Dao, 2023), (Zhu et al., 2024).
+Specifically, in the SSM module, Mamba utilizes parameters ($\Delta$, $A$, $B$, $C$) to transform an input sequence $x(t)$ into an output sequence $y(t)$ using a hidden state $h(t)$.
 The discretization process converts $A$ and $B$ into $\bar{A}$ and $\bar{B}$, respectively, using the time step size parameter $\Delta$.
 Structured state space models, inspired by continuous systems, can be computed similarly to RNNs or in the form of global convolution due to their linear time invariance (LTI) property.
-Mamba introduces a selective property to structured state space model, tying parameters to the current input, which breaks the LTI property and hinders parallel training.
-To address this, Mamba employs a hardware-aware algorithm, enabling its SSM module to possess the selective property and perform parallel training.
+Mamba introduces a selective property to the structured state space model, tying parameters to the current input.
+This breaks the LTI property and hinders parallel training.
+To address this, Mamba employs a hardware-aware algorithm, enabling its SSM module to possess the selective property while performing parallel training.
 To be specific, the discretization process can be expressed as:
-$\bar{A}$ $=exp(\Delta A)$ (22)
-$\bar{B}$ $=(\Delta A)^{-1}(exp(\Delta A)-I)\cdot\Delta B$
+$\bar{A}$ $=\exp(\Delta A)$ (6)
+$\bar{B}$ $=(\Delta A)^{-1}(\exp(\Delta A)-I)\cdot\Delta B$
 After that, the calculation in Mamba can be expressed as:
@@
 Delta projection is a module in SSM that’s learnable and tasked with transforming the parameter $\Delta$.
-Since adapting the delta projection alone can indirectly adapt the entire SSM module (i.e., $\bar{A}$ and $\bar{B}$ are all determined by $\Delta$), it is the most critical component of the SSM module.
-We investigate the performance of two adaptation strategies: adapting only the input linear projection layer (denoted as "in") and adapting both the input linear projection layer and SSM (denoted as "$\Delta$ + in" since we only adapt delta projection in SSM module).
+Since adapting the delta projection alone can indirectly adapt the entire SSM module (i.e., $\bar{A}$ and $\bar{B}$ are determined by $\Delta$), it is the most critical component of the SSM module.
+We investigate the performance of two adaptation strategies: adapting only the input linear projection layer (denoted as "in") and adapting both the input linear projection layer and the SSM (denoted as "$\Delta$ + in" since we only adapt the delta projection in the SSM module).
 Our results, as shown in Table 1, indicate that applying Lily solely to the input projection yields better performance than applying it to both the input and delta projection modules.
@@
 Additionally, developing PEFT methods specifically tailored to Mamba-based models, whether for vision or language foundation models, is also a promising direction for future work.
-Appendix F Performance with Different Learning Rates
+Appendix G Performance with Different Learning Rates
 Since we only tuned the learning rate in the commonsense reasoning experiment, we provide the performance of commonsense reasoning under different learning rates in Table 16.
@@
 3e-4 72.9 85.6 77.8 92.7 83.3 89.7 77.6 82.8 82.8
-Appendix G Does Selectivity Help?
+Appendix H Does Selectivity Help?
 Figure 8: Implementation of Lily with no selectivity.
@@
 As shown in Table 17, removing selectivity from Lily results in generally poorer performance compared to vanilla Lily.
-This is likely because the lack of selectivity causes Lily to simply aggregate all the HP expert, leading to inferior performance.
-This validates the design choice of using routers in Lily to selectively allocate weights to HP experts, rather than simply summing them.
+This is likely because the lack of selectivity causes Lily to simply aggregate all the $B$ expert, leading to inferior performance.
+This validates the design choice of using routers in Lily to selectively allocate weights to $B$ experts, rather than simply summing them.
 Table 17: Commonsense reasoning results of Lily without selectivity. We provide results using two learning rates.
@@
 5e-4 71.3 85.5 78.1 84.3 79.6 86.4 76.1 79.0 79.8
-Appendix H How to Allocate Parameters?
-Since Lily alters the traditional LoRA’s layer-bound setup, increasing the parameters of Lily can be achieved through two approaches: 1) increasing ne, i.e., increasing the number of LP and HP experts, and 2) increasing the rank, i.e., increasing the parameter size of each individual LP or HP expert.
+Appendix I How to Allocate Parameters?
+Since Lily alters the traditional LoRA’s layer-bound setup, increasing the parameters of Lily can be achieved through two approaches: 1) increasing ne, i.e., increasing the number of $A$ and $B$ experts, and 2) increasing the rank, i.e., increasing the parameter size of each individual $A$ or $B$ expert.
 In this section, we investigate which factor has the greatest impact on performance.
@@
 The results are shown in Fig.
-9, from which we observe that more LP and HP experts with smaller rank (i.e., bigger ne and smaller rank) generally performs worse.
+9, from which we observe that more $A$ and $B$ experts with smaller rank (i.e., bigger ne and smaller rank) generally performs worse.
 We argue that this is because, although increasing the attention granularity allows for finer details, the resulting performance gain is not as significant as the gain obtained by increasing the rank, i.e., increasing the model’s capacity to learn more information.
@@
 The hyperparameter ne is specifically tuned to maintain the same amount of parameter count for a fair comparison.
-Appendix I More on Subject-driven Generation
+Appendix J More on Subject-driven Generation
 We provide more results on subject-driven generation in Fig. 10 and Fig. 11.
@@
 Figure 11: More subject-driven generation results for subjects that are reported in the experiment section.
-Appendix J More on Attention Maps of Lily and LoRA
+Appendix K From a Feature Merging Perspective
+Apart from having higher-rank weight updates than LoRA, Lily also enables comprehensive information access across layers.
+Lily enables access to information or features from all other layers when adapting a target module at a specific layer thanks to the inter-connectivity of the adapters.
+We aim to understand how Lily achieves this comprehensive information access from the perspective of visual tasks as shown in Fig.
+12.
+We can observe that, in Lily, the distinctness of the attention maps between layers is not as pronounced as in LoRA.
+This validates Lily’s ability to enable all-level information access, since adaptation at each layer takes into account features from other layers.
+Additionally, we specifically visualize the actual feature differences between different layers in Fig.
+13.
+We observe that Lily has more points with low feature differences (blue color) than LoRA, indicating that the distinctness of features between layers in Lily is generally lower than in LoRA.
+This further demonstrates Lily’s ability to enable comprehensive information access.
+Although we enable all-level information access, what prevents the features from becoming completely identical is the selectivity introduced by Lily, which we specify in the following section.
+Figure 12: Attention maps of Lily and LoRA.
+The input images for the example here are taken from Caltech101 datasets from VTAB-1K benchmark.
+It can be observed that features from a certain layer have more similarity to those in other layers in Lily than in LoRA.
+Figure 13: Feature difference measured in absolute distance for each element.
+We compare Lily and LoRA in terms of the difference between features from different layers.
+In this example image taken from Caltech101, we visualize the feature difference between layers 6 and 1, as well as between layers 6 and 9.
+Appendix L More on Attention Maps of Lily and LoRA
 We provide more visualization results of the attention map from both LoRA and Lily on Caltech101 dataset from VTAB-1K benchmark in Fig. 14.